### v1

# One More Glance with Sharp Eyes: Rethinking Lightweight Captioning as a Practical Visual Specialist

Abstract
Image captioning is fundamental for applications like video instruction systems and exploration robots, yet deploying such models on local devices is challenging due to the high computational demands of multimodal large language models (MLLMs). To address this, we first explore lightweight captioning by implementing a specialist based on a 125M-parameter language model, 56 times smaller than LLaMA-7B, and evaluating its performance on both single-sentence and detailed captioning tasks. Surprisingly, we find that our model can achieve performance comparable to large multimodal generalists, suggesting its potential to serve as a strong visual specialist for on-device applications. While promising, our model also exhibits a limitation: like other MLLMs, it suffers from visual blindness, occasionally resulting in semantic captioning errors. We carry out toy experiments and investigate the underlying causes, where we observe that the problems arise from ineffective attention mechanisms and limited visual representations. To alleviate them, we develop a novel captioning framework, Sharp-Eyed Refinement, which enhances caption quality through improved visual grounding. At its core, our DeepLens extracts detailed visual representations by concentrating on informative regions identified during the initial glance. Our experiments confirm both the advantages of our specialist over prior small captioning models and large generalists and the effectiveness of our framework.
Figure 1: (a) Image captioning serves as the cornerstone of diverse applications, such as robots for assistive navigation. (b) The natural process of image description is illustrated: humans first take in the overall scene, then glance at specific regions to notice finer details. (c) Our Sharp-Eyed Refinement framework mimics the human tendency to take a second, more focused glance, allowing the captioning specialist to revise and improve the initial descriptions.
1 Introduction
Image captioning has rapidly advanced with the emergence of MLLMs [43; 37], leveraging the impressive language processing capabilities of LLMs [94]. This captioning technology can play a crucial role in various applications, as shown in Fig. 1 (a). For example, video instruction systems utilize frame-wise caption generation for temporal understanding [76; 83; 78], while exploration robots construct graph-structured scene descriptions to navigate complex environments [82; 28].
Despite these advantages, deploying such technologies on local devices remains a significant challenge due to the large computational demands of MLLMs [29]. Hence, we explore lightweight captioning by implementing a captioning specialist using OPT-125M [92], a model 56 times smaller than LLaMA-7B [67] in the LLaVA framework [43]. Evaluating it on both single-sentence and detailed captioning tasks, we found that our lightweight model can achieve impressive performance. In fact, it not only surpasses recently introduced small-scale captioning models [58] but also achieves performance comparable to large-scale multimodal generalists, such as LLaVA-7B [43] and Instruct-BLIP [15]. These results are particularly notable considering the limited language modeling capacity of OPT-125M, and suggest that the complex reasoning capabilities of LLMs may be less critical for image captioning tasks centered on factual information.
Despite these remarkable results, our lightweight captioner is not without limitations. It occasionally generates inaccurate captions, an issue widely studied in MLLMs, such as LLM’s hallucination [5; 2] and the blindness of vision encoders [65; 66]. Since the majority of our specialist’s parameters are dedicated to the visual encoder, we focus on visual blindness as a potential root cause. This issue implies that since the vision encoder manipulates visual representation while perceiving the image with its eyes wide shut [66], the LLM often struggles to generate accurate outputs.
We take a step towards analyzing the limitations of current captioning models. First, we inspect the model’s attention during a single pass captioning and find that the model struggled to focus on critical regions of the image, instead spreading its attention across the entire image. Moreover, we observe that visual representations from a CLIP encoder [56] lack sufficient detail for accurate caption generation. Specifically, when reconstructing images using these features, significant deviations from the original images occurred. These analyses highlight the need for more focused and detailed visual representations to provide the LM with clear detection information.
Building on these insights, we propose a novel captioning framework, Sharp-Eyed Refinement, as illustrated in Fig. 1. Our method mirrors the human process of image description, where an initial broad understanding of the image is followed by a more focused and detailed analysis. In our framework, our model generates a rough caption, then refines its description. At the heart of our framework is our multimodal connector, DeepLens, which facilitates efficient re-examination of both the image and the initial caption during refinement.
Our contributions are summarized as follows: (i) We investigate the practically important yet underexplored topic of lightweight captioning models (Sec. 2), and demonstrate their potential as strong visual specialists (Sec. 3). (ii) We suggest embodying the human mechanism of ‘one more glance with sharp eyes’ and introduce a novel refinement framework (Sec. 4). (iii) We present DeepLens, which extracts relevant visual features guided by the initial caption to enable focused refinement (Sec. 5). (iv) We conduct extensive experiments to validate the effectiveness and efficiency of our proposed framework and module. (Sec. 6). It is our hope that these efforts will inspire further development of efficient captioning models deployable on resource-constrained devices.
Figure 2: We evaluate our lightweight captioner on both single-sentence and detailed captioning tasks. It uses 93% fewer parameters and offers 97% faster inference than LLaVA-1.5-7B.
2 Motivation & Scope
Image Captioning as Foundational Technology
Image captioning is the process of transforming visual information from an image into a natural language description [51; 24]. More than just a standalone task, this technology serves as a fundamental component for a wide range of applications, as outlined in Fig. 1 (a). To process video input in chatbot systems [61; 77; 95; 83], image captioning is employed to generate descriptions of multiple frames, which are then integrated into LLMs as prompts to facilitate instruction following. Similarly, exploration robots [82; 28] designed for disaster environments, as well as assistive robots for the visually impaired, utilize captioning to encode observed scenes into structured graph representations, supporting navigation and user interactions. In this work, we focus on image captioning as a crucial enabler for these applications.
Real-World Deployment Challenges.
Recent studies aiming to implement such applications commonly utilize open-source MLLMs [76] or cloud-based APIs (e.g., OpenAI API) [82] as image captioners. From a practical application perspective, however, these approaches present significant challenges, as (1) open-source MLLMs require substantial computational resources that exceed the capabilities of edge devices like smartphones [7; 63], while (2) cloud-based APIs depend on a stable internet connection, which is likely to be unavailable in disaster situations. Tab. 1 presents the number of parameters used by recent MLLMs and their NVIDIA GPU memory requirements, assuming FP16 precision. Given that these applications require repetitive captioning process across multiple scenes [90], the feasibility of current models for real-world deployment remains uncertain. This resource-demand mismatch highlights the need for lightweight captioning models.
Edge devices LLaVA-1.5 7B LLaVA-NeXT 34B LLaVA-OA 72B
iPhone 16 Galaxy S25 mPLUG-Owl3 8B InternVL 40B Qwen2-VL 72B
8GB 12GB 16GB ++ 68GB ++ 140GB ++
Table 1: Available memory on edge devices and GPU usage of recent MLLMs by parameter size.
3 Exploring Lightweight Captioning
3.1 Model Construction & Experimental Setup
Model construction. To construct a lightweight captioning model, we initially focused on minimizing reliance on LLMs, as they contribute to the majority of computational overhead in MLLMs (e.g., 96% in LLaVA-7B [43] is attributed to LLaMA [67]). To this end, we simply build our captioning specialist by replacing the LLaMA-7B used in LLaVA construction with OPT-125M [92], a language model 56 times smaller. Released in 2022, OPT-125M is significantly compact unlike modern small-scale language models exceeding 500M parameters.
Implementation details. We leverage the publicly available LLaVA-1.5 [42] code. Other than this modification, all training strategies follow the original code, including batch size, learning rate, and updating both the language model and the connector during fine-tuning while freezing the vision encoder. More details are available in Sec. H.1 and at https://github.com/junha1125/Lightweight-Captioner. All our experiments are conducted using two NVIDIA A6000 GPUs.
Datasets and Evaluations. The multimodal connector is first pretrained on the Caption Concept-balanced 558K [42], after which the model is fine-tuned on task-specific datasets such as COCO [14], DCI [70], or ShareGPT4V [11], as introduced later. We consider traditional metrics such as BLEU [55], CIDEr [71], and BERTScore [93], as well as MLLM-as-a-Judge [9; 8] with GPT-4o-mini.
3.2 Results on Single Sentence Captioning
MS COCO Captions [14], the most widely used captioning benchmark, consists of 113k images and 5 captions per image. One caption typically contains a single sentence of about 10 words. The results in Tab. 2 show that our model outperforms previous small captioning works [25; 72] that utilize fewer than 1B parameters. Specifically, our model achieves CIDEr scores that are 6.9 higher than SmallCap [58]. Considering that our model employs a small LM and does not involve any specialized methods, these results may represent surprisingly strong performance. Furthermore, we compare our specialist model against large multimodal generalists [13; 3] that leverage LLMs and are trained on proprietary instruction-following datasets [43]. Despite requiring significantly fewer computational resources, our specialist model achieves comparable performance to these larger counterparts. A discussion of these results is provided in Tab. 14 and Sec. D.1.
MS COCO [14]
model venue # data # params B@4 [55] MET [16] CIDEr [71] BERT [93] CLAIR [8] GPT [9]
G InstructBLIP [15] NeurIPS23 130M> 8.2B 38.0 29.4 127.8 69.1 - -
Unified-IOXL [46] ICLR23 130M 7.3B 37.0 29.5 123.6 68.2 - -
Shikra [10] arXiv23 - 7.2B - - 117.5 - - -
[…]
ShareGPT4V [11] & DCI [70]
model venue # data # params B@4 [55] CIDEr [71] BERT [93] CAPT [18] CLAIR [8] GPT [9]
G InstructBLIP [15] NeurIPS23 130M> 8.2B - 28.2 - 41.7 - -
[…]
SmallCap* [58] CVPR23 0.5M 450M 14.5 20.1 28.9 23.3 - -
[…]
GLaMM [60]
model venue # data # params B@4 [55] CIDEr [71] BERT [93] CAPT [18] CLAIR [8] GPT [9]
[…]
Table 2: We evaluate our specialist and existing captioning models. Here, G represents a large multimodal generalist trained with instruction data to handle multiple tasks, while "S" denotes the small-scale captioning specialist trained on a dedicated captioning training dataset. * indicates models that are fine-tuned in this study, as they were not trained for detailed captioning tasks.
3.3 Results on Detailed Captioning
Images can be described with more detailed content beyond single-sentence descriptions [21]. Therefore, we continue our study by examining our specialist’s capabilities on more challenging detailed captioning tasks. For this experiment, we perform fine-tuning with the DCI [70] and ShareGPT4V [11] datasets, containing 100K and 7.4K training images, respectively. We then evaluate our model on the test sets of the above datasets. We also use the GLaMM [60] dataset, which contains 100K training and 6K testing images. Dataset details are included in Sec. F.2.
Figure 3: Generation examples from the original OPT-125M. Despite such limited capabilities, lightweight language models can perform factual image captioning efficiently.
We initially anticipated that our specialist would struggle with this challenging task, given that OPT-125M possesses limited language processing capabilities, presumably due to its small parameter count, as shown in Tab. 3. However, the results in Tab. 2 and Fig. 4 reveal that our model performs surprisingly well on detailed captioning tasks, contrary to our expectations. We note that we keep the vision encoder frozen while only updating OPT-125M. Thus, detailed caption generation relies entirely on OPT-125M, with the vision encoder providing only visual cues.
Figure 4: Qualitative results on detailed captioning. Our specialist, despite its small size, can generate effective image descriptions. More examples are provided in Sec. E.
These results underscore the efficiency of a lightweight LM for captioning tasks. Although MLLMs rely on the complex reasoning abilities of LLMs for tasks such as VQA and instruction following, our experiments indicate that captioning, primarily involving the enumeration of factual objects, does not significantly require such capabilities of LLM. This suggests that a lightweight captioning model can serve as an efficient and practical alternative for captioning-based applications.
Note 1:
Factual image captioning does not significantly require the complex reasoning abilities of LLMs and can be handled by a lightweight language model.
Figure 5: Visualization of attention maps for red-highlighted words: (left) when the model generates the caption in a single pass, and (right) when it performs one more glance with the initial caption.
Figure 6: Reconstructed images from CLIP visual encoder embeddings. We compare results using last-layer features (LF) and multi-layer features (MF). For MF, we extract representations from layers 13, 18, and 23 of the 24-layer ViT-L encoder.
Despite its impressive performance, our specialist shares similar limitations with other MLLMs [65; 45], occasionally producing incorrect predictions. Prior works have interpreted these errors from two perspectives: (i) the hallucination hypothesis, where LLMs, during the reasoning process, describe non-existent elements in the image [5; 39]; (ii) the visual blindness hypothesis, where the visual encoder generates image embeddings which are unclear or ambiguous [66]. Given that our model primarily consists of parameters dedicated to the vision encoder (e.g., 67% of total parameters reside in this component), we hypothesize that the observed errors are more likely a result of visual blindness rather than hallucination. In the following section, we provide an in-depth analysis of this issue and introduce our proposed solution.
4 Why Sharp-Eyed Refinement?
In this section, we analyze the limitations of our captioning specialist by proposing two hypotheses and validating them through toy experiments and visualizations. We then introduce our captioning framework to alleviate these issues.
Analysis 1.
(Hypothesis.) We propose a first hypothesis: they fail to focus on essential critical regions (e.g., entities, attributes, relations), instead maintaining a redundant and blurry view of the entire image. This may stem from the current captioning framework, which requires the model to observe the image and generate the entire caption in a single pass. This approach can be likened to asking a human to describe an image in a single glance, without focusing on specific details. (Verification.) To investigate this issue, we analyze which regions the model attends to during the generation of specific words. As detailed in Sec. H.3, the attention maps are visualized by implementing code that references techniques from the API [88]. As illustrated in Fig. 5, the model often exhibits diffuse, global attention and fails to focus on specific regions. These findings underscore the necessity for an approach that guides the captioning process toward more effective attention mechanisms.
Note 2:
Single-pass captioning can lead specialists to adopt broad attention, neglecting critical visual regions necessary for precise description.
Analysis 2.
(Hypothesis.) Eyes Wide Shut [66] shows that visual modality deficiencies-specifically, the CLIP vision encoder [56] yielding similar representations for distinct images (a phenomenon termed visual blindness)-undermine MLLM performance. Building on these findings, we hypothesize that the widely used CLIP features may produce ambiguous representations. (Verification.) We attempt to reconstruct images from their feature embeddings. The quality of the reconstructed images serves as an indicator of how much visual detail the features preserve and, consequently, their effectiveness in supporting captioning tasks. For this experiment, we directly employ the decoder architecture from Masked Autoencoder (MAE) [22]. We keep the CLIP encoder frozen while training only the MAE decoder on the COCO images. Our goal was to visualize how the image is represented within the language model. As illustrated in Fig. 6, the images reconstructed from CLIP embeddings exhibit significant deviations from the original images and appear less clear. This result indicates that such visual representations may not provide sufficiently detailed visual information for captioning.
Note 3:
CLIP’s final-layer features provide only coarse-grained visual information, limiting their suitability for fine-grained captioning tasks.
One More Glance.
Building on the above insights, we propose a novel captioning framework, Sharp-Eyed Refinement. As illustrated in Fig. 1 (b) and (c), our framework is inspired by the human description process. Specifically, the model first generates an initial caption. Then, the same LM, in conjunction with our proposed multimodal connector DeepLens, refines this output. This refinement step mimics a human taking a second glance to revise their initial impression, allowing the model to generate captions with more effective attention mechanisms and fine-grained grounding (i.e., with sharp eyes). The design and functionality of DeepLens are described in the following section.
Our refinement process is similar to ideas explored in LLM research through self-refinement [50; 32; 54], where models improve their outputs by revisiting previous generations. Extending this concept, we apply refinement to captioning models, uniquely integrating visual features for multimodal refinement. To the best of our knowledge, this represents the first attempt of its kind.
5 Inside Our Framework: DeepLens and Learning Strategy
We aim to mitigate the above limitations by incorporating our proposed DeepLens into the caption refinement process. In particular, unlike prior multimodal connectors [43; 37; 3], DeepLens takes two new inputs: (i) previously generated captions and (ii) multi-layer features from the vision encoder. Below, we examine the significance of this design in detail.
Looks at what matters.
Imagine receiving an image along with the caption “A cat relaxing on a brown chair” and being instructed to refine it. In such a scenario, humans naturally focuses on key elements such as the cat and the chair. We attempt to employ a similar mechanism by providing the previously generated caption as an additional input to guide the refinement process. This design enables the module to identify and attend to the relevant visual regions. Similar to the experiment in Sec. 4, we also inspect the attention maps while the model takes a second glance using the previous captions, as shown in Fig. 5. We observe that the model focuses more precisely on regions corresponding to specific words, indicating that it attempts to look at what matters in the image.
Looks in detail.
To mitigate the coarse-grained visual feature problem, some studies proposed using additional vision towers [65; 66; 27], but this increases computational costs. For example, Interleaved MoF [66] adds 300M parameters for DINOv2 [53]—a substantial 60% increase relative to our 500M parameter model. Instead, we focus on maximizing the utility of the existing vision encoder. We design DeepLens to leverage features extracted from various layers of the vision encoder, as these multi-level features enable more semantically rich encoding of detailed image elements, following FPN [41]. Indeed, the reconstruction experiment conducted in Fig. 6 demonstrates that utilizing these features helps generate clearer images. This implies that the visual embeddings manipulated by DeepLens using multi-level features contain sufficient fine-grained information about the image, supporting the LM in recognizing visual content more accurately.
Note 4:
Encouraging attention to key regions via the initial caption, along with multi-level utilization of a single vision encoder, can be an effective and efficient strategy.
The architectural design of DeepLens follows a BERT [17] structure that takes two inputs: visual features and projected embeddings of the initial captions. Further details are in Sec. C.
Figure 7: Qualitative results of Our Sharp-Eyed Refinement. More examples are in Sec. E.
Figure 8: To develop our specialist, our model undergoes a 2-step fine-tuning process for initial caption generation and refinement. During the inference stage, the caption is initially generated and then refined, following the green and blue lines, respectively.
Fine-tuning Strategy.
Assume we have a training set $X=\{x_{1},\dots,x_{N}\}$, where each $x_{k}=(i_{k},c_{k})$ consists of an image $i_{k}$ and its corresponding ground-truth caption $c_{k}$; $N$ denotes the total number of such pairs. As illustrated in Fig. 8, we adopt a two-stage fine-tuning strategy. In the first stage, we focus on enabling the model to generate an initial caption by following the training procedure of LLaVA. Specifically, we perform supervised fine-tuning on the MLP connector and the lightweight language model using this dataset.
The second stage aims to facilitate the model’s ability to learn the refinement process effectively. We first consider that directly using the momentary generated captions as initial inputs and the true captions $c_{k}$ as the targets for refinement can be misleading. Suppose our captioning specialist were to initially produce something like ‘a brown table in front of a window’ while the true caption is ‘a cat sitting on a chair’; these two descriptions share little in common and provide no meaningful signal for gradual correction. Training on such disparate pairs would encourage the model to ignore the provided initial caption and generate a wholly new one, rather than learn to refine it. Additional explanations are given in Sec. D.5.
To address this challenge, we construct pseudo-initial captions $\hat{c}_{k}$ that remain close to their corresponding ground-truth $c_{k}$. Specifically, we leverage GPT-4o-mini [1] to possibly introduce small errors into each $c_{k}$, producing slightly altered versions that focus on entities, attributes and relations (e.g., ‘a dog sitting on a chair’). During second-stage fine-tuning, each pair $(i_{k},\hat{c}_{k})$ is fed into DeepLens alongside the language model, and a cross-entropy loss is applied. Our fine-tuning regimen consists of 10 epochs for the first stage and 2 epochs for the second stage. The details and examples of how $\hat{c}_{k}$ is generated are provided in Sec. F.1. The resulting dataset $X$, including the pseudo-initial captions $\hat{c}_{k}$, will be made publicly available.
Rationale for Refinement Fine-tuning.
Let $\pi_{\theta}$ denote the language model. Each pseudo-initial caption $\hat{c}_{k}$ deviates from the ground truth $c_{k}$ at only a few token positions $E_{k}=\{t\mid\hat{c}_{k,t}\neq c_{k,t}\}$. Under the sequence-level cross-entropy objective
$\mathcal{L}(\theta)\;=\;-\!\mathbb{E}\Bigl{[}\sum_{j}\log\pi_{\theta}\!\bigl{(}c_{k,j}\mid i_{k},\hat{c}_{k},c_{k,<j}\bigr{)}\Bigr{]},$ (1)
gradients are likely to be primarily concentrated on tokens in $E_{k}$, leading to a form of targeted optimization, in which the model retains the correct parts of $\hat{c}_{k}$ while rewriting only the erroneous ones. Writing $\Delta_{k}(\theta)=\log\pi_{\theta}(c_{k}\mid i_{k},\hat{c}_{k})-\log\pi_{\theta}(\hat{c}_{k}\mid i_{k},\hat{c}_{k}),$ we have $\mathcal{L}(\theta)\propto-\mathbb{E}[\Delta_{k}]$; hence, minimizing $\mathcal{L}$ is equivalent to maximizing the expected margin $\Delta_{k}$, thereby directly increasing the likelihood of the refined caption relative to its flawed precursor. In effect, each gradient step encourages DeepLens to focus on visual regions likely responsible for errors in the initial caption and guides the language model to better interpret these refined features, resulting in more accurate captions. We note that this targeted optimization resembles the philosophy of Direct Preference Optimization (DPO) [57], if we consider $c_{k}$ and $\hat{c}_{k}$ as preferred and less-preferred responses. In contrast to DPO, which treats both responses symmetrically during optimization, our method offers a new perspective by assigning them distinct roles as input and target.
Inference Strategy.
Given an image $i_{k}$ and instruction $p$, the model first generates an initial caption $o_{initial}$, storing intermediate ViT [19] features in a buffer $B$ to avoid redundant computation. $o_{initial}$ is then fed into DeepLens and the LM to produce the refined output $o_{refined}$. Although this process can be carried out iteratively, similar to Self-Refine [50] in LLM research, our experiments indicate that it does not provide additional performance gains. We therefore do not adopt a repeated process but anticipate it to be a promising direction for future research, as described in Sec. D.6 and Sec. B.6.
MS COCO [14]
model #params B@4 [55] gain CIDEr [71] gain CLAIR [8] gain GPT [9] gain
[…]
+ SeR with 1 500M 39.6 +0.2 131.9 +2.3 - - - -
+ SeR with 2 500M 39.6 +0.2 132.3 +2.7 - - - -
Single glance with 2 450M 39.7 +0.3 130.6 +1.0 - - - -
ShareGPT4V [11] & DCI [70]
model #params CIDEr [71] gain CAPT [18] gain CLAIR [8] gain GPT [9] gain
[…]
GLaMM [60]
model #params CIDEr [71] gain CAPT [18] gain CLAIR [8] gain GPT [9] gain
[…]
Table 3: Quantitative evaluation of Sharp-Eyed Refinement (SeR). The results demonstrate the effectiveness of SeR in improving caption quality. We also present an ablation study on the key inputs to DeepLens: ① initial captions and ② multi-layer features.
6 Empirical Validation
model time
LLaVA 274.49 s
[…]
Figure 9: Inference time required to generate captions for 100 streaming images.
Ablation Study of DeepLens. To assess the contributions of DeepLens, we conduct an ablation study as shown in Tab. 3 and Fig. 7. The results show that our framework yields both quantitative and qualitative improvements in caption quality. The refinement process enhances initial captions by +3.9 and +2.5 CIDEr points for single-sentence and detailed captioning tasks, respectively. While the integration of multi-layer visual features (i.e., with ②) provides modest gains, its effectiveness remains constrained by the limitations of single-pass generation. In fact, our method introduces additional parameters (approximately 50M for DeepLens) and an additional inference step with OPT-125M. However, considering the significant advantages shown in Tab. 9, we believe that the computational overhead introduced by our refinement method is minimal. This measurement was conducted over 100 iterations with a batch size of 1.
Evaluation on Long Range Video Question Answering.
Since captioners are expected to serve as foundational components in downstream applications as mentioned in Sec. 2, it is important to verify their utility in practical settings. To this end, we evaluate various captioners on the Long-Range Video Question Answering task introduced in [90]. Following the baseline setup of Zhang et al. [90], each captioner generates captions for multiple frames from videos spanning several minutes to hours. These captions are then fed into an LLM to answer video-related questions. As shown in Tab. 4, our specialist achieves competitive accuracy with significantly fewer parameters and faster inference than large-scale generalists, demonstrating its potential for real-world applications discussed in Fig. 1. The reported time includes both the captioning and LLM inference steps, assuming a 10-minute video processed at the frame level. The small specialists [58; 25], including ours, use weights fine-tuned on the ShareGPT [11] and DCI [70] datasets.
LLM Captioner LLaVA-1.0 [43] BLIP-2 [37] LLavA-1.5 [42] SmallCap* [58] Tag2text* [25] Our specialist + SeR
Qwen2.5 14B [84] #params 7.3B 7.4B 7.3B 450M 900M 450M 500M
accuracy 47.6 50.6 51.1 41.8 47.1 49.3 50.8
time 29m 20s 29m 44s 29m 20s 4m 45s 7m 14s 4m 53s 5m 10s
Table 4: Evaluation on Long-Range VideoQA. We follow the baseline setup of LVQA [90], replacing the captioner with our specialist and large multimodal generalists. Captions from all frames are aggregated and fed into Qwen2.5-14B to perform the final video question answering.
Note 5:
Despite its significance, the use of lightweight captioners has been underexplored. Our proposed baseline and framework demonstrate the potential of lightweight captioning for on-device applications.
7 Related Work
Multimodal Large Language Models (MLLMs)
have attracted considerable research attention due to their versatile applications, such as chat-bots [80]. Early approaches integrated contrastive image-language pretrained models [74] with powerful LLMs, enabling complex reasoning. The development of instruction-based datasets [43] and innovative training strategies [35; 86] have further accelerated progress, substantially improving MLLM performance and broadening their capabilities. Despite these achievements, most MLLMs heavily depend on large-parameter LLMs, making deployment on memory-constrained devices unfeasible. This limitation will likely restrict access for a significant portion of users worldwide.
Image Captioning Models.
Recent advances in image captioning have improved training efficiency and descriptive fidelity. Approaches such as CaMEL [6] and SmallCap [58] emphasize minimizing trainable parameters by leveraging mean-teacher distillation and employing retrieval augmentation, while Tag2Text [25] and LoCCa [72] introduce novel mechanisms such as dedicated tagging and location-aware refinement to improve caption quality. Unlike existing approaches that focus on reducing trainable parameters, or rely on single-pass inference—potentially missing crucial details—our method prioritizes inference efficiency considering on-device operation and systematically addressing the limitations of single-pass generation.
Visual Blindness in VLMs.
Despite significant advancements, MLLMs still face limitations in their visual capabilities, hindering their practical applications. Eyes Wide Shut [66] demonstrated that even GPT-4V [1] struggles with basic visual questions. Research on this topic typically points to two main sources of failure: one relates to the language decoder, which can hallucinate details not present in the image [5], while the other focuses on the visual encoder, which may provide ambiguous visual information. Several studies, including Cambrian [65], believe the visual encoder is a critical bottleneck. We also concentrate on the visual issue, particularly within the context of our lightweight model, where the visual encoder accounts for a significant portion of the parameters. Furthermore, we introduce a novel operational framework to improve visual grounding.
Self-Refinement in LLMs.
Humans often refine their writing through iterative review to enhance clarity and precision [50; 20]. Recent research has applied this refinement concept to LLMs, introducing techniques. For instance, Self-Refine [32; 59; 54] enabled models to autonomously critique and iteratively enhance their outputs. Unlike such approaches limited to the LLM domain, our method introduces refinement in a multimodal context, guided by both language and vision.
8 Conclusion & Broader Impact
[…]
Broader Impact.
We demonstrate the potential of small-scale specialists, highlighting their surprising effectiveness. As a next step, future research may explore iterative refinement strategies, integrating external tools (e.g., zoom or crop, as in GPT-4o [52]), and designing a unified multimodal connector for both the initial and secondary glance. More research questions that aim to enhance both captioning performance and efficiency in real-world applications are provided in Sec. D.6.
Acknowledgments
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (RS-2022-00143911,AI Excellence Global Innovative Leader Education Program)
Appendix
[…]
Appendix A Summary of Our Work
Motivation
  1. 1.
    Image captioning serves as the cornerstone of diverse applications such as assistive devices.
  2. 2.
    However, running multimodal large language models for this task is challenging for edge devices.
  3. 3.
    We highlight the practical importance of lightweight captioning.
What We Have Demonstrated?
  1. 1.
    Factual image captioning does not significantly require the complex reasoning abilities of LLMs and can be handled by a lightweight language model.
  2. 2.
    Single-pass captioning can lead specialists to adopt broad attention, neglecting critical visual regions necessary.
  3. 3.
    CLIP’s final-layer features provide only coarse-grained visual information, limiting their suitability for fine-grained captioning tasks.
Novel Caption Refinement Framework
  • What is it? Humans first take in the overall scene, then glance at specific regions to notice finer details. Our Sharp-Eyed Refinement framework mimics this human tendency, allowing the captioning specialist to revise and improve initial descriptions.
  • Our novel insight: [Uncaptioned image] Encouraging attention to key regions via the initial caption, along with multi-level utilization of a single vision encoder, can be an effective and efficient strategy.
  • Why novel? To the best of our knowledge, this is the first work to apply multimodal refinement that jointly utilizes visual features and the model’s initial textual output.
    • $\circ$
      NLP community: Humans often refine their writing, and coders revise their code through iterative review. Recent research has applied this refinement/correctness concept to LLMs [50, 59, 81, 31]. Unlike such approaches limited to the LLM domain, our method introduces refinement in a multimodal context.
    • $\circ$
      Multimodal LLMs community: Existing models [43, 3, 96] heavily rely on the complex reasoning capabilities of LLMs and typically generate outputs in a single pass. We uniquely demonstrate the effectiveness of revisiting and refining its own textual outputs.
    • $\circ$
      Visual Blindness community: Several studies [65, 66] believe the visual encoder is a critical bottleneck and adapting multiple vision encoders. However, this requires additional parameters. Since our research focuses on a lightweight model, we maximize the utility of the existing vision encoder through multi-level utilization.
    • $\circ$
      Lightweight captioning community: This field has been gradually declining since the remarkable capabilities of LLMs were discovered. In this work, we revisit the practically important yet underexplored topic of lightweight captioning models.
        Unlike ours, some methods [48, 58, 73] utilize prior captions obtained through a heavy image-text retrieval process. For example, SmallCap performs similarity matching between a given image and 500,000 candidate captions at each iteration.
        Unlike ours, several works [25, 47] incorporate object detection or tagging procedures. However, despite these extra components, their captioning performance has been limited.
        Unlike ours, which refines text output generated via a full forward pass, certain works [48, 12] adopt Diffusion Transformers and denoise latent text embeddings. Moreover, they provide little motivation or analysis as to why such a process is important from the perspective of utilizing visual cues more effectively.
Broad Comparisons
  1. 1.
    This is the first paper to demonstrate the amazing capability of a lightweight captioning model beyond single-sentence captioning, including not only detailed captioning but also long-form VideoQA.
  2. 2.
    Unlike traditional works, we compared not only existing small specialist models but also large multimodal models, demonstrating their potential as strong visual specialists.
Positive Influences
  1. 1.
    Industry: Our model can be easily applied in industrial scenarios such as assistance devices for the visually impaired or navigating robots. Our model runs successfully on actual edge devices like Jetson Nano.
  2. 2.
    Academy: For researchers with limited access to large-scale GPUs who are unable to operate LLMs, our factual image captioning model offers a promising baseline.
Appendix B Additional Experiments
B.1 When GPT meets ‘one more glance’
In this work, we suggest that allowing the model to take one more glance, guided by the initial caption and aligned visual features, offers an effective approach, as discussed in Note 4. We further explore how OpenAI’s GPT models, which have demonstrated remarkable performance, might be influenced by our refinement framework. While our framework consists of two components, taking one more glance and refining with sharp eyes through DeepLens, we apply only the former as modifying their internal architecture is not feasible. The results are shown in Fig. 10, where GPT-4-turbo is used. As a commercial model with a sufficiently large number of parameters, it often produces an accurate initial caption without noticeable errors, but we still observed occasional inaccuracies in its outputs. To explore whether refinement could help, we instructed the model to take one more glance and found that GPT can generate a refined caption by re-examining the given image. OpenAI’s GPT-o3 reflects an awareness of this idea through the introduction of the Thinking with Images [52]. However, this concept remains less explored in the research community, suggesting the need for further investigation.
B.2 When larger specialists meet ‘one more glance with sharp eyes’
Building on our lightweight specialist, we naturally ask: since our framework and module offer a scalable methodology, can they also be effective when applied to larger multimodal models? To investigate this, we scaled the language model from OPT-125M to OPT-1.3B and also utilized LLaMA-2-7B. Specifically, we fine-tuned these models on the ShareGPT and DCI training datasets to serve as captioning specialists (i.e., single pass). Next, we incorporated our Sharp-Eyed Refinement (i.e., SeR) framework with DeepLens and further fine-tuned them using the datasets described in Sec. 5. As shown in Tab. 5, our framework consistently improves caption accuracy across different models. In particular, we observe gains of 1.2 and 0.9 points in CAPT [18] evaluation for OPT-1.3B and LLaMA-2-7B, respectively. Additionally, the results indicate that using larger language models to build captioning specialists leads to improved performance, observed in prior LLM scaling studies [30, 79, 24]. Despite the gains observed with larger models such as OPT-1.3B and LLaMA-2-7B, we note that their considerable size (roughly 10x and 56x larger than OPT-125M, respectively) may hinder real-world applicability. It is therefore crucial to continue exploring captioning systems that balance effectiveness with efficiency.
ShareGPT4V [11] & DCI [70]
vision total CIDEr [71] CAPT [18]
Language model encoder LoRA [23] # params single pass + SeR single pass + SeR
OPT-1.3B [92] CLIP ViT-L $\times$ 1.6B 50.2 53.1 (+2.9) 49.0 50.2 (+1.2)
LLaMA-2-7B [67] CLIP ViT-L 7.3B 57.3 61.7 (+4.4) 52.7 53.6 (+0.9)
Table 5: Captioning performance of larger specialists with and without Sharp-Eyed Refinement (SeR). SeR consistently improves accuracy across model scales, including OPT-1.3B and LLaMA-2-7B.
Figure 10: When GPT meets ‘one more glance’. Example outputs from OpenAI’s GPT before and after being prompted to take one more glance. The refined captions illustrate its potential responsiveness to visual re-examination.
B.3 Evaluation on an Actual Edge Device
To highlight the practical value of our approach, we present an evaluation of our lightweight captioner on edge devices, where large-scale MLLMs are typically infeasible to deploy. Our core motivation is rooted in the observation that while MLLMs are powerful, they are computationally intensive and difficult to run in resource-constrained environments. In contrast, lightweight captioners remain underexplored despite their potential for real-world applications.
Here, we demonstrate that lightweight captioners can be effectively deployed on edge devices such as RTX 3090, and Jetson Nano. We evaluate captioning performance on MS COCO and ShareGPT4V&DCI datasets, and additionally measure inference time, memory usage, and power consumption. All models were executed for 100 iterations with a batch size of 1. The results in Tab. 6 show that our method consistently performs well across devices, and notably, remains fully operational on setups where models like LLaVA-1.5 fail to run. These findings support the deployability of our framework in edge scenarios and encourage further exploration of lightweight captioning for real-world assistive technologies.
B.4 Comparison on different learning strategies
As writers frequently revise their drafts to improve clarity or tone, and programmers refactor code to enhance structure, we demonstrate that allowing multimodal models to refine their own outputs through our framework can be an effective strategy. In this subsection, we examine how our lightweight specialist performs under different learning strategies. The first strategy involves training with as much data as possible. Here, we train the model on a combined set of COCO, ShareGPT, DCI, and GLaMM to maximize data exposure (i.e., more data). The second strategy explores distillation using teacher-generated captions. Specifically, we train our model on captions produced by LLaVA-1.6-34B [44] (i.e., distillation). Our original strategy was straightforwardly fine-tuning the model exclusively on task-specific data (i.e., origin). The results are summarized in Tab. 7. For the ‘more data’ strategy, we observe limited performance, likely due to reduced task alignment caused by mixing heterogeneous datasets. In contrast, the ‘distillation’ strategy demonstrates that learning from a strong teacher can improve performance even in a single-pass setting. Notably, when combined with our refinement framework (SeR), it yields further gains, suggesting that distillation and SeR can be effectively integrated. Additionally, as noted in Sec. 5, the datasets we used share similarities with those in Direct Preference Optimization (DPO). Investigating the impact of training our model using reinforcement learning methods such as DPO could be a promising direction for future work.
Resource (RAM) Data Model Mem. Inf. time Power. B@4 [55] CIDEr [71] CLAIR [8] GPT [9]
Jetson Nano (4G) All LLaVA-1.5-7B out-of-memory N/A N/A N/A N/A N/A N/A
[…]
Table 6: Evaluation results across different hardware resources and datasets.
ShareGPT4V [11] & DCI [70]
Metric origin +SeR gain more data distillation +SeR gain
CIDEr [71] 40.5 43.6 +3.1 36.8 42.6 43.6 +2.0
CAPT [18] 45.9 48.4 +2.5 43.2 46.5 47.4 +0.9
Table 7: Performance comparison of our lightweight captioner under different learning strategies: origin, more data, and distillation. Distillation from a strong teacher, especially when combined with Sharp-Eyed Refinement (SeR), leads to the best results.
B.5 Efficacy of Multi-Level Features with Other Vision Encoders
We investigate whether our strategy of selecting multi-level features can generalize to vision encoders beyond CLIP [56], which is our original setup. As shown in Tab. 8, the approach consistently improves performance across multiple encoders. Specifically, both SigLIPv2 [68] and DINOv2 [53] exhibit performance gains when equipped with multi-level features, demonstrating that our method is not limited to CLIP-based models. Interestingly, CLIP with multi-level features outperforms the combination of CLIP and DINOv2 in several evaluation metrics, while maintaining greater parameter efficiency. On the other hand, DINOv2 alone yields relatively lower performance, likely due to weaker alignment with language features. For these experiments, we pair OPT-125M with each vision encoder and fine-tune the resulting models on the ShareGPT [11]&DCI [70] datasets. Subsequently, we apply our Sharp-Eyed Refinement framework using the selected multi-level features and perform second fine-tuning with the datasets described in Sec. 5 of the main paper.
Vision encoder #params indices of selected layers CIDEr [71] CLAIR [8] GPT [9]
CLIP [56] 300M {23} 42.8 55.8 2.78
300M {13, 18, 23} 43.3 57.7 3.02
CLIP [56]+DINOv2 [53] 600M {23} + {23} 42.9 57.3 3.02
SigLIPv2 [68] 300M {23} 43.0 56.2 2.88
300M {15, 23} 45.9 57.7 3.05
300M {13, 18, 23} 45.5 58.2 3.07
DINOv2 [53] 300M {23} 32.8 48.6 2.55
300M {13, 18, 23} 33.0 50.1 2.66
Table 8: For the captioning specialist, we evaluate different vision encoders and the corresponding selected layer indices.
B.6 Iterative Sharp-Eyed Refinement
To explore the potential of iterative refinement beyond a single pass, we conduct additional experiments using captioning specialists with larger model capacity. This decision is motivated by our initial hypothesis that the constrained capacity of small models may have hindered the effectiveness of multi-step refinement. As shown in Tab. 9, applying multiple refinement iterations to the smallest language model, OPT-125M, results in only limited improvements. In contrast, we observe that multi-step refinement leads to performance gains when used with higher-capacity captioners. These findings underscore the potential of scaling up refinement strategies, as similarly observed in recent LLM-based studies [50, 59, 81, 31]. We plan to further investigate this direction in future work.
For implementation, we construct larger captioning specialists based on two configurations: LLaMA-2-7B + CLIP-ViT-Large and OPT-1.3B + CLIP-ViT-Large. These models are initially fine-tuned on the ShareGPT and DCI training datasets via single-pass training. We then apply the Sharp-Eyed Refinement (SeR) framework with DeepLens in multiple iterations, followed by additional fine-tuning using the datasets described in Sec. 5.
Language model vision encoder Benchmarks Eval. matrix initial cap. refinement refinement *2 refinement *3 refinement *4
OPT-125M CLIP ViT-L MS COCO CIDEr [71] 129.6 133.5 133.8 133.1 133.4
OPT-1.3B 133.7 135.7 138.1 138.1 138.5
LLaMA-2-7B 137.4 141.4 141.6 141.9 141.7
[…]
Table 9: Multiple iterative refinements with larger language models. This experiment demonstrates better refinement capability as model capacity increases.
Appendix C DeepLens
C.1 Architectural details
We have shown that our refinement framework can improve caption quality. At the core of it lies DeepLens, which we hypothesize plays a critical role by aligning multi-level visual features with semantic cues from the initial caption. As illustrated in Fig. 11, DeepLens takes two new inputs: ① the previously generated caption, represented as token embeddings, and ② multi-layer features extracted from the vision encoder. Let $N$, $d$, and $m$ denote the number of visual tokens, the embedding dimension, and the number of selected layers, respectively. The selected features are concatenated along the channel dimension to retain hierarchical visual information. These fused inputs are then passed through Transformer layers following the BERT architecture [17]. The resulting output is provided to the language model, enabling refinement conditioned on a focused and semantically grounded visual representation.
Figure 11: Illustration of the DeepLens architecture, which integrates ① token embeddings from the initial caption and ② multi-layer visual features for refinement.
C.2 Ablation study
We conduct a series of ablation experiments to analyze the design choices behind DeepLens, as summarized in Tab. 10. The corresponding model variants are illustrated in Fig. 12. Our study spans three key aspects: connector selection, refinement configurations, and the layer indexing strategy for visual feature extraction. (I) We begin by comparing different multimodal connectors used in the initial caption generation phase. While prior works have explored a range of designs, such as Cross-Attention [58, 6], Q-Former [38], and Transformer-style modules, we find that the simple MLP connector from LLaVA [42] offers competitive performance. Notably, (b) and (f), which adopt multi-level features and a BERT-style Transformer block respectively, perform slightly better than the MLP-based connector. However, given the marginal gains, we retain the lightweight MLP structure for efficiency. (II) Next, we evaluate several connectors for the refinement phase (SeR). Among them, combining our base configuration (a) with structure (A), which utilizes both proposed inputs, yields strong performance. Although we also applied (A) to configuration (b), the improvement was marginal, so we did not adopt it as our baseline. (III) Finally, we investigate the impact of selecting different sets of layers from the ViT encoder. We test multiple combinations and find that using a diverse set of layers {13, 18, 23} results in the highest performance.
Figure 12: Architectural designs of the model variants used in the ablation study. Each configuration corresponds to a specific component or setting evaluated in Tab. 10
(I) connectors for initial generation
(a) (b) (c) (d) (e) (f)
129.6 ✓ 130.6 125.9 127.7 122.9 130.9
(II) (a) + connectors for SeR (b) + con.
(A) (B) (C) (D) (A)
133.5 ✓ 131.9 132.1 131.9 133.6
(III) indexes of selected layers in ViT
{23} {13, 23} {15, 23} {15, 19, 23} {13, 18, 23}
131.9 133.0 133.0 133.3 133.5 ✓
Table 10: Ablation results evaluating connector types, refinement configurations, and ViT layer selections in DeepLens. Architectural variants are illustrated in Fig. 12.
Appendix D Discussions
D.1 Why previous small models fall short?
The development of captioning models can be divided into two eras: before and after the adoption of LLMs. Prior to the integration of LLMs, many approaches relied on architectures that did not involve a significant number of parameters. For example, CaMEL [6] and SmallCap [58] employed GPT-2 models with 125M to 350M parameters, while Tag2Text [25] and LoCCa [72] used BERT models ranging from 300M to 900M parameters. The emergence of models such as ClipCap [51], BLIP [37], and LLaVA [43] shifted the direction of the captioning field. With growing interest in the impressive capabilities of LLMs, much of the research has focused on building MLLMs.
In contrast to this trend, we analyze small captioning models and uncover the following: Aforementioned small captioning models typically adopt an architecture where visual features are injected via cross-attention. Unfortunately, when using a small language model, this design does not appear to be particularly beneficial. Empirically, this limitation is reflected in the performance of structure (c) in Fig. 12, which yields a relatively low score of 125.9 as shown in Tab. 10. Had the field not been so heavily influenced by the rise of LLMs, such limitations in small model architectures might have been identified and addressed earlier.
We adopt the LLaVA architecture and inject visual features into the self-attention inputs, as shown in Fig. 12 (a). We show that this simple setup surprisingly delivers stronger performance (i.e., Note 1 in Sec. 3) and contributes to demonstrating that small models can be viable for practical applications. We hope this encourages a shift toward reducing dependence on LLMs in specific tasks like captioning, and sparks further interest in the development of lightweight yet effective models.
D.2 Limitations of Existing Multimodal Large Language Models
We examine the broad challenges faced by existing multimodal large language models (MLLMs), particularly in addressing the issue of visual blindness. Prior works such as Eyes Wide Shut [66] and Cambrian [65] have identified and attempted to mitigate this issue through the use of multiple vision encoders, including DINOv2 [53], SigLIPv2 [68], and CLIP [56]. However, as illustrated in Fig. 13, even models with large parameter sizes continue to exhibit difficulty in generating consistent long-form captions in complex, multi-object scenes. We also conduct experiments using two recent capable MLLMs, LLaVA-Next [36] and LLaVA-OneVision [34]. Despite incorporating the sophisticated approach which is to split an image into grids and utilize all the features from each grid, both models continue to suffer from incorrect captioning results.
These findings suggest that visual blindness is a persistent limitation across model scales and architectures. In this context, we believe that our Sharp-Eyed Refinement framework—which encourages attention to key regions via initial captions and leverages multi-level features from a single vision encoder—as a meaningful and efficient step toward addressing this challenge.
D.3 Limitations of lightweight captioners
While we have demonstrated the potential of small models in captioning tasks, the use of lightweight LMs unavoidably introduces some limitations. In particular, we observe that the model occasionally suffers from issues such as repetitive phrasing, reduced fluency, limited OCR capability, and a lack of general world knowledge. Examples of these cases are provided in Fig. 14. These limitations may stem from two primary factors: (i) the small number of parameters, which can restrict the model’s capacity for complex reasoning and language generation [30], and (ii) the limited scale and quality of training data, as our model was trained on approximately 500K image-caption pairs from ShareGPT-4V [11] which contains machine-generated captions. A natural direction for future work is to investigate how far the capabilities of small models can be scaled with access to larger and higher-quality training datasets. In addition to the results in Sec. D.2, we observe similar issues in larger models such as LLaVA-1.5, suggesting that these challenges remain unresolved [5, 66] and require deeper investigation.
D.4 Limitations of existing evaluation methods
To ensure fair comparison across captioning models, we adopt seven evaluation metrics: BLEU@4 [55], METEOR [16], CIDEr [71], BERTScore [93], CAPTURE [18], CLAIR [8], and MLLM-as-judge [9] (i.e., GPT-based evaluation). For CLAIR and MLLM-as-judge, we randomly select 100 image samples and evaluate them using 10 different seeds to compute both mean and standard deviation. The values reported alongside the $\pm$ symbol indicate the standard deviation across seeds. Despite these efforts, we find that current evaluation methods do not always align well with human judgment. For instance, certain models perform better under specific metrics but worse under others, leading to inconsistent rankings. Moreover, small specialists including ours sometimes exhibit deficiencies in fluency, as discussed in Sec. D.3, which existing metrics struggle to capture. As we examined, MLLM-as-judge yields relatively reliable results. In contrast, CLAIR showed higher variance across runs, limiting its consistency.
These observations highlight the need for further research into robust captioning evaluation. In particular, future work should account for multiple dimensions of quality, including fluency, coherence, faithfulness, relevance, informativeness, and completeness. Moreover, moving beyond traditional n-gram-based matching, reliable evaluation methods powered by MLLMs (e.g., OpenAI GPT) are needed. We believe such advancements can support greater trust in captioning-based applications.
Figure 13: Despite their size, existing MLLMs struggle with visual blindness, especially in complex scenes. We hope our approach offers a meaningful step toward alleviating this issue.
Figure 14: Despite their efficiency, small models suffer from various limitations. Further research is required to assess how high-quality data can compensate for their weaknesses.
D.5 Impact and Ablation of Pseudo-Initial Captions in Stage-2 Fine-tuning
We first provide a discussion, with examples, on how the inclusion of pseudo-initial captions leads to better supervision during training. We highlight the following points: (i) The model is trained to generate the ground-truth (GT) caption given both the image and the pseudo-initial caption as input: $I+\hat{c}\;\rightarrow\;\text{Our model}\;\rightarrow\;c$. (ii) If the pseudo-initial caption is generated following the strategies in Sec. F.1, then it is unlikely to include too many differing tokens from the GT caption.
Suppose we have “A woman in a room with a cat.” as the $c$ and $\hat{c}$ in Tab. 11.
pseudo-initial caption $\hat{c}$ different tokens $E={t\mid\hat{c}(t)\neq c(t)}$ in Sec. 5
A woman in a room with two dogs two / dogs
A cat sitting on a chair in front of the window. sitting / on a chair / in front of the window
Table 11: Token-level differences between pseudo-initial and ground-truth captions.
In such cases, if each pair of pseudo-initial caption and GT caption contains a reasonable number of overlapping tokens, the model is encouraged to refer to the pseudo-initial caption and selectively revise the differences to produce the GT caption. Conversely, if the pair contains too many differing tokens, the model is likely to avoid relying on the pseudo-initial caption and instead generate the GT caption independently. This mechanism helps the model refine incorrect parts rather than regenerate the caption from scratch, thereby mitigating misleading supervision.
As an additional experiment, we fine-tune our captioner via Stage 2 as in Fig. 8 using pseudo-initial captions generated under the following four conditions: (Data 1) Initial captions generated by the captioner fine-tuned via Stage 1, (Data 2) Pseudo-initial captions that are significantly different from the GT caption, (Data 3) Pseudo-initial captions with minor modifications from the GT caption, which is our original strategy, and (Data 4) Pseudo-initial captions consisting of (i) one generated as in Data3, and (ii) the GT caption itself (e.g., GT: “A woman in a room with a cat”; pseudo-initials: “A boy in a room with a dog” and “A woman in a room with a cat”).
Detailed captioning CIDEr [71] gain CAPT [18] gain GPT [9] gain
[…]
Table 12: Detailed captioning results on ShareGPT4V [11] & DCI [70] with fine-tuning on four types of pseudo-initial captions. The comparison highlights how different levels of overlap or inclusion of GT captions affect model performance.
Tab. 12 demonstrates that Data 1 provides reasonable gains within our framework, while Data 2 leads to the anticipated issue where the model tends to ignore the initial caption and generate a new one. Interestingly, Data 4 performs comparably to Data 3, and together with the results from Data 1, indicates that our framework is robust to moderate variation in pseudo-initial caption quality.
D.6 Further Research Questions
To guide future exploration, we outline several research directions that may advance the field of lightweight captioning and multimodal learning more broadly:
  1. 1.
    What is the minimal model size required for a captioning specialist to be practically useful in real-world assistive technologies? Where does the performance-efficiency trade-off stabilize?
  2. 2.
    Although supervised fine-tuning is commonly used in multimodal training, recent work has shown the benefits of reinforcement learning methods in the LLM field [67]. Can similar reward-based training be applied to enhance captioning models?
  3. 3.
    The LLaVA framework remains a popular baseline due to its simplicity and open-source accessibility. However, could a unified architecture, such as a fully integrated Transformer-based multimodal model, serve as a better foundation for building practical captioners?
  4. 4.
    Multimodal models heavily rely on the quality of the representations produced by vision encoders. To what extent can we trust visual features from models like SigLIP [89] or MAE [91], especially for fine-grained captioning tasks?
  5. 5.
    Similar to self-refinement in LLM research [50], how can we enable our framework to iteratively improve its own outputs?
  6. 6.
    Can our framework be effective in simple VQA settings? Is an LLM necessary in such setups?
  7. 7.
    Do potential biases (e.g., gender, occupation, and age) exist in captioning models, and if so, how can we address them together with cultural contexts and deployment domains such as indoor, outdoor, and robotic environments? Could domain adaptation [64] with user feedback [62] serve as a solution?
We consider these questions promising directions for future research.
Figure 15: Additional results corresponding to the analyses in Fig. 5 and Fig. 6. It illustrates diffuse attention patterns (top) and the limited visual detail captured by CLIP features (bottom).
Appendix E Additional Experimental Results
Analysis of captioning operation.
We present additional results from the analysis experiments conducted in Sec. 3, as shown in Fig. 15. These results support our earlier findings that single-pass captioners tend to adopt diffuse attention patterns, and that CLIP’s final-layer features may lack the granularity needed for fine-grained description.
Comparison between our specialist and LLaVA-1.5-7B.
Additional results comparing our lightweight specialist with the large multimodal generalist (LLaVA-1.5-7B), as discussed in Sec. 3, are shown in Fig. 16. Despite its limited parameter size and a frozen vision encoder, our specialist demonstrates unexpectedly strong performance on captioning benchmarks, indicating that small models may still handle detailed captioning tasks reasonably.
Impact of Sharp-Eyed Refinement.
Additional results related to our refinement framework are presented in Fig. 17, building on the findings discussed in Sec. 6. Our refinement method yields caption-quality improvements in both single-sentence and detailed captioning.
Appendix F Extensive Details on The Dataset
F.1 How to generate the pseudo-initial caption $\hat{c}_{k}$?
We describe our strategy to generate the pseudo-initial caption as the following. The instruction prompt used for this strategy is shown in Fig. 19.
  • Pseudo-initial captions are generated by modifying 0–3 points from the ground-truth (GT) caption.
  • Modifications are based on four categories: Entity (e.g., chair, cat), Attribute (e.g., color, material like wooden, count like three cups, texture, shape, size, inspired by DSG [1]), Relation (e.g., A in front of B, or B in front of A), Action (e.g., eating, blowing).
  • Except the modifications, the overall sentence structure and style should be preserved.
  • It can occasionally be identical to the GT caption.
  • Pseudo-initial captions are generated by modifying 0–3 points from the ground-truth (GT) caption.
  • Few-shot examples are provided like the examples in Tab. 13.
GT caption pseudo-initial caption
A woman in a room with a cat A woman in a room with two cats
A woman in a room with a dog
A woman in a room with a cat
A multicolored motorcycle rests outside of a sheep farm A multicolored bicycle rests inside a sheep farm
A bright red motorcycle rests outside of a sheep farm
A multicolored motorcycle races around a sheep farm
Table 13: Few shot examples of GT captions and corresponding pseudo-initial captions.
We consistently generate three pseudo-initial captions per GT caption. Hence, the dataset can be summarized as Tab. 14, where ‘#pairs’ indicates the number of image-caption pairs.
Dataset #images #GT captions per image #pairs for fine-tuning stage 1 #pseudo-initial captions per GT caption #pairs for fine-tuning stage 2
COCO [14] 113K 5 565K 3 1.6M
ShareGPT4V [11] 100K 5 500K 3 1.5M
DCI [70] 7.4K 10 74K 3 0.2M
GLaMM [60] 550K 1 550K 3 1.6M
Table 14: Statistics of datasets used for fine-tuning. Notably, the number of images in fine-tuning stages 1 and 2 are identical.
Figure 16: Qualitative comparison between our lightweight specialist and the large multimodal generalist LLaVA-1.5-7B using MS-COCO [14] (top) and ShareGPT4V [11] & DCI [70] (bottom). Despite its smaller size and simpler architecture, our model produces competitive descriptions.
Figure 17: Qualitative examples of our sharp-eyed refinement. The results show improved caption quality after refinement, using MS-COCO [14] (top) and ShareGPT4V [11] and DCI [70] (bottom).
F.2 Datasets for Detailed Captioning
While MS COCO [14] has long served as the standard benchmark for image captioning, its annotations are typically limited to single-sentence descriptions that often fail to capture the richness of visual content. Recent works [87, 21, 40, 18] have highlighted this limitation and emphasized the need for more detailed captioning datasets. In this study, we leverage three datasets that offer relatively high-quality and more comprehensive image descriptions: ShareGPT-4V [11], DCI [70], and GLaMM [60]. For ShareGPT-4V, captions are initially generated by GPT-4o and subsequently refined by human annotators. DCI contains human-authored captions. Finally, GLaMM generates detailed captions by aggregating outputs from numerous open-source tools, including object detectors and scene parsers, and composing them using an LLM.
F.3 ShareGPT & DCI
The DCI dataset [70] contains 7.4K training images and 0.4K test images, with each image paired with 10 human-written captions averaging 55 words in length.
The ShareGPT4V dataset [11] includes 100K images, each accompanied by a single long and human-verified caption of approximately 200 words. Using this format directly poses a challenge for fair evaluation with n-gram-based metrics, which benefit from having multiple reference captions per image. So, we summarized the original caption into five captions, each containing roughly 50 words, since generating multiple captions while preserving the original length could introduce hallucinations. The prompt used for this summarization process is presented in the next section, and the resulting dataset will be made publicly available.
Given the relatively small scale of DCI compared to MS COCO (which contains 118K images with five captions per image), we combined DCI with the processed ShareGPT4V dataset to construct a unified benchmark. This yielded 102.4K training images (100K from ShareGPT4V and 7.4K from DCI), each with ten or five detailed captions, and 5K test images—comparable in scale to MS COCO.
F.4 GLaMM
The GLaMM dataset [60] contains automatically generated captions produced using a combination of object detection, scene-graph parsing models, and LLMs. On average, each caption consists of approximately 45 words. In our experience, despite GLaMM’s reliance on extensive visual tools, the quality of the generated captions was often inconsistent. Frequent factual inaccuracies were observed, for example, one to two incorrect words appearing in every two and three captions. Each image in GLaMM is paired with a single caption, which presents challenges for n-gram-based evaluation, as previously discussed. Nevertheless, we include this dataset in our experiments. We randomly sampled 600K image-caption pairs from the full dataset, allocating 30K for testing and using the remaining 570K for training. Possibly due to the quality issues mentioned above, models trained on GLaMM generally underperformed compared to those trained on ShareGPT4V and DCI.
Appendix G Prompt Templates
We share the prompt templates used when interacting with OpenAI’s GPT models throughout our study. First, the prompt used for summarizing long captions into shorter ones, introduced in Sec. F.3, is shown in Fig. 18. Second, the prompt used to generate pseudo initial captions for our refinement framework, as described in Sec. 5, is provided in Fig. 19. The examples generated using it are in Fig. 21. Lastly, the prompt used in the MLLM-as-judge evaluation setup is shown in Fig. 20. We note that we strictly follow the template proposed in the original MLLM-as-judge papers [8, 9].
Figure 18: Prompt used for summarizing long-form captions into shorter, multi-reference captions.
Figure 19: Prompt used to generate pseudo initial captions for Sharp-Eyed Refinement.
Figure 20: Prompt used for MLLM-as-judge evaluation, following the original template from [8, 9].
Figure 21: Examples of pseudo initial captions generated using the prompt in Fig. 19, from MS COCO (top) and ShareGPT4V & DCI (bottom).
Appendix H Experimental Details
H.1 Pretraining and Finetuning
Our implementation is initialized from the LLaVA-1.5 repository [42]. To reduce reliance on LLMs, we replace LLaMA with the OPT series of language models [92]. The training process consists of three stages: pretraining, finetuning for caption generation (in Sec. 3), and finetuning for sharp-eyed refinement (in Sec. 5). The hyperparameters used for each stage are summarized in Tab. 15. We closely follow the original LLaVA training setup. As emphasized in Sec. D.1, our method differs from previous small-model approaches in that the visual features are directly used as self-attention inputs to the language model.
Pretraining Fine-tuning
Dataset LCS-558K [42] MS COCO [14] or SharedGPT [11] + DCI [70]
Adapter 4-layer MLP with GELU 4-layer MLP with GELU / Deeplens
Trainable Adapter layers only Adapters + Language Model
Training Epochs 1 10 / 2
Learning Rate $1\times 10^{-4}$ $2\times 10^{-5}$
Weight Decay 0 0
Warm-up Ratio 0.03 0.03
Learning Rate Scheduler Cosine decay Cosine decay
Table 15: Hyperparameters used for model training. The settings to the left and right of the / correspond to those used in Sec. 4 and Sec. 3, respectively.
H.2 Experiment setting of large multimodal generalist
Tab. 2 compares specialist models with several generalist MLLMs, including InstructBLIP [15], Unified-IO-XL [46], Shikra [10], Qwen-VL [3], and LLaVA-1.5 [42]. These generalist models were reported to be trained on MS COCO images, but not on datasets such as ShareGPT4V, DCI, or GLaMM. Instead, they were instruction-tuned on large-scale datasets; for example, Qwen-VL and InstructBLIP were trained with approximately 1.5B and 130M instruction samples, respectively. For this reason, we categorize them as generalist models.
The generalist MLLMs were evaluated using publicly available checkpoints from their official GitHub repositories, without any additional fine-tuning. For the single-sentence captioning task, we used the prompt: “Provide a one-sentence caption for the provided image.” For the detailed captioning task, we used the prompt: “Describe the photo within 55 words.”. Additionally, we conduct an experiment on LLaVA-1.5 with the alternative instruction, as shown in Tab. 16. These results reveal two important points: (1) evaluation metrics such as BERTScore [93] penalize long-form outputs, and (2) longer generations tend to induce more hallucinations, consistent with prior observations [26].
LLaVA-1.5 CIDEr [71] BERTScore [93] CAPT [18]
“Describe the photo within 55 words” 36.1 36.6 48.6
“Describe the photo in detail” 12.8 17.6 40.6
Table 16: Performance of LLaVA-1.5 under different prompt instructions.
H.3 Attention map visualization
As part of our analysis in Sec. 4, we evaluate whether the model attends to the appropriate image regions when generating specific words in a caption. To this end, we adapt the visualization code provided by API [88], originally designed to point out attention maps between images and questions in VQA tasks. We modify the attention hooking module (https://github.com/yu-rp/apiprompting/blob/master/API/API_LLaVA/hook.py) to visualize attention between image regions and selected words within captions. This allows us to inspect which parts of the image the LM focuses on when producing certain tokens, such as the phrase ‘a white toilet’ in Fig. 5.
H.4 Image reconstruction
To investigate whether CLIP’s visual representations are coarse or ambiguous, we conducted an image reconstruction experiment using a Masked AutoEncoder (MAE) framework in Sec. 4. In this setup, a ViT encoder produces visual features, which the decoder then uses to reconstruct the image. We adopt the same visual encoder as used in LLaVA-1.5, ‘CLIP ViT-L/14-336’, and keep its parameters frozen. The decoder receives the visual embeddings without masking and predicts the corresponding RGB image. We utilize two types of visual inputs: (i) last-layer features (i.e., LF), and (ii) multi-level features (i.e., ML), where ML consists of outputs from layers 13, 18, and 23 of the encoder. The decoder is trained on 100K images from the MS COCO dataset. Further architectural details are available in the MAE repository, https://github.com/facebookresearch/mae/blob/main/models_mae.py. The hyperparameters for optimizing the decoder are summarized in the table below.
Optimizer Learning Rate $\beta$ Weight Decay Batch Size Epoch
AdamW 1e-5 (0.9, 0.95) 0.05 64 10
Table 17: Hyperparameters used for training the MAE decoder in the image reconstruction experiment on MS COCO.
H.5 Long Range Video Question Answering
To assess the feasibility of deploying lightweight captioners in real-world applications, we evaluate our model on the Long-Range Video Question Answering task [90]. This task requires the system to answer multiple-choice questions based on user queries about videos that are 10 minutes or longer in duration. We emphasize that there are currently no video MLLMs capable of handling this task directly. Most existing models [33, 4, 96] impose limits on the number of input video frames they can process, making it difficult to cover the full temporal span of long videos.
To address this, recent approaches [85, 75] propose first extracting per-frame captions and then injecting both the captions and the question into an LLM, which can handle over 100K tokens in a single prompt. For evaluation, we follow the setup provided in the official implementation of LLoVi (https://github.com/CeeZh/LLoVi), replacing only the captioning models. When using the generalist model, we generate a caption for each frame using the prompt: "Describe this frame within 50 words". For our specialist model, we use the same prompt and apply the version fine-tuned on the ShareGPT4V [11] and DCI [70] datasets.