### v2

# One More Glance with Sharp Eyes: Rethinking Lightweight Captioning as a Practical Visual Specialist

Abstract
Image captioning is fundamental for applications like video-grounded chatbot systems and navigation robots, yet deploying such models on local devices is challenging due to the high computational demands of multimodal LLMs (MLLMs). To address this, we first build lightweight captioning models using a 125M-parameter language model, 56 times smaller than LLaMA-7B, and evaluate their performance not only on single-sentence but on detailed captioning tasks. We obtain surprising results showing that our model can achieve performance comparable to MLLMs, suggesting its potential to serve as a strong captioning specialist for on-device applications. While promising, our model also exhibits a limitation: like other MLLMs, it suffers from occasional captioning errors. We investigate the underlying causes and observe that the problems stem from ineffective attention mechanisms and limited visual representations. To alleviate them, we develop a novel captioning framework, Sharp-Eyed Refinement, which enhances caption quality by refining coarse descriptions into more precise captions. At its core, DeepLens improves visual grounding by re-examining the informative regions identified in the initial glance. Experimental results demonstrate the superiority of our model over both recent lightweight captioning methods and MLLMs in detailed captioning and even in long-range video QA tasks.
Figure 1: (a) Image captioning serves as the cornerstone of diverse applications, such as assistive navigation systems. However, existing captioning models are often too computationally heavy for deployment or suffer from limited performance. (b) The natural process of image description is illustrated: humans first take in the overall scene, then glance at specific regions to notice finer details. (c) Our Sharp-Eyed Refinement framework mimics the human tendency to take a second, more focused glance, allowing the captioning specialist to revise and improve its captioning outputs.
1 Introduction
Recent progress in image captioning has been driven by the emergence of MLLMs llava ; li2023blip2 , which exploit the advanced language processing capabilities of LLMs zhao2023survey . Building on these advances, image captioning has become a crucial component in various applications, as shown in Figure 1 (a). For example, video-based chatbot systems utilize frame-wise caption generation for temporal understanding wang2023internvid ; moment-gpt ; videotree , while navigation robots construct graph-structured scene descriptions to operate in complex environments xie2024embodied ; jiang2024roboexp .
Despite recent progress, deploying such technologies on local devices remains a significant challenge due to the large computational demands of MLLMs jin2024efficient . Motivated by this, we explore lightweight captioning by implementing a captioning model using OPT-125M zhang2022opt , a model 56 times smaller than LLaMA-7B touvron2023llama in the LLaVA framework llava . To assess its effectiveness, we evaluate it on both single-sentence and detailed captioning tasks, and find that our lightweight model achieves impressive performance. In fact, it not only surpasses previous lightweight captioning models ramos2023smallcap ; huang2023tag2text ; vipcap but also achieves performance comparable to MLLMs, such as LLaVA-7B llava and Instruct-BLIP dai2023instructblip . Taken together, these findings reveal that the complex reasoning capabilities of LLMs are less critical for image captioning tasks that primarily involve enumerating factual details.
Despite these remarkable results, our lightweight model is not without limitations. It occasionally generates inaccurate captions, an issue widely studied in MLLMs, such as LLM’s hallucination bai2024hallucination ; an2024agla and the blindness of vision encoders cambrian ; eyeswideshut . Since the majority of our model’s parameters are dedicated to the visual encoder, we focus on visual blindness as a potential root cause. This issue means that since the vision encoder manipulates visual representation while perceiving the image with its eyes wide shut eyeswideshut , the LLM often struggles to generate accurate outputs.
We take a step towards analyzing the limitations of current captioning models. First, we inspect the model’s attention during a single pass captioning and find that the model struggles to focus on critical regions of the image, instead spreading its attention across the entire image. Moreover, visual representations from a vision encoder like CLIP CLIP lack sufficient detail for accurate caption generation. Specifically, when reconstructing images using these features, significant deviations from the original images are observed. These results highlight the need for more focused and detailed visual representations to provide the LM with clear detection information.
To address these limitations, we propose a novel captioning framework, Sharp-Eyed Refinement, as illustrated in Figure 1. Our method mirrors the human process of image description, where an initial broad understanding of the image is followed by a more focused and detailed analysis. In our framework, our model generates a rough caption, then refines its description. At the heart of our framework is our multimodal connector, DeepLens, which facilitates efficient re-examination of both the image and the initial caption during refinement. The effectiveness of our framework is demonstrated through experiments on single-sentence, detailed captioning, and long-range video QA; representative results are summarized in Figure 2.
Our contributions are summarized as follows: (i) We investigate the practically important yet underexplored topic of lightweight captioning (Section 2). We study lightweight models and demonstrate their potential as strong captioning baselines (Section 3). (ii) Inspired by the human mechanism of ‘one more glance with sharp eyes,’ we introduce a novel refinement framework (Section 4, 5). (iii) We conduct extensive experiments to validate the effectiveness and efficiency of our proposed framework (Section 6). We hope these efforts will inspire further development of efficient captioning models.
Figure 2: We evaluate our lightweight captioner on both single-sentence and detailed captioning tasks. It uses 93% fewer parameters and offers 97% faster inference than LLaVA-1.5-7B.
2 Motivation & Scope
Image Captioning as Foundational Technology.
Image captioning is the process of transforming visual information from an image into a natural language description mokady2021clipcap ; hu2022scaling . More than just a standalone task, this technology serves as a fundamental component for a wide range of applications, as outlined in Figure 1 (a). To process video-grounded chatbot systems sarto2023positive ; wang2024internvideo2 ; zhao2023learning ; moment-gpt , image captioning is employed to generate descriptions of multiple frames, which are then integrated into LLMs as prompts to facilitate instruction following. Similarly, exploration robots xie2024embodied ; jiang2024roboexp designed for disaster environments, as well as assistive robots for the visually impaired, utilize captioning to encode observed scenes into a structured graph, supporting navigation and user interactions. In this work, we focus on image captioning as a crucial enabler for these applications.
Real-World Deployment Challenges.
Recent studies aiming to implement such applications commonly utilize open-source MLLMs wang2023internvid or cloud-based APIs (e.g., OpenAI API) xie2024embodied as image captioners. From a practical perspective, however, these approaches present significant challenges, as (1) open-source MLLMs require substantial computational resources that exceed the capabilities of edge devices like smartphones cai2020tinytl ; song2023ecotta , while (2) cloud-based APIs depend on a stable internet connection, which is likely to be unavailable in disaster situations. Table 1 presents the number of parameters used by recent MLLMs and their NVIDIA GPU memory requirements, assuming FP16 precision. Given that these applications require repetitive captioning process across multiple scenes zhang2023simple , the feasibility of current models for real-world deployment remains uncertain. This resource-demand mismatch highlights the need for lightweight captioning models.
Edge devices LLaVA-1.5 7B LLaVA-NeXT 34B LLaVA-OA 72B
iPhone 16 Galaxy S25 mPLUG-Owl3 8B InternVL 40B Qwen2-VL 72B
8GB 12GB 16GB ++ 68GB ++ 140GB ++
Table 1: Available memory on edge devices and GPU usage of recent MLLMs by parameter size.
Generalist vs. Specialist.
Following previous works ye2024mplug ; mome ; wu2024visionllm , we use the term generalists to denote models trained on diverse data for broad purposes (e.g., current MLLMs llava15 ; llavanext ), and specialists to denote models trained solely on task-specific data (e.g., small captioning models ramos2023smallcap ; vipcap )."
3 Exploring Lightweight Captioning
To address the challenges outlined above, this section presents our design of a lightweight captioning specialist and provides comprehensive analyses of its effectiveness.
Model construction. To construct a lightweight captioning model, we initially focus on minimizing reliance on LLMs, as they contribute to the majority of computational overhead in MLLMs (e.g., 96% in LLaVA-7B llava is attributed to LLaMA touvron2023llama ). To this end, we build our captioning specialist by replacing the LLaMA-7B used in LLaVA construction with OPT-125M zhang2022opt , a language model 56 times smaller. It is notable that OPT-125M is significantly compact unlike modern small-scale language models exceeding 500M parameters.
Implementation details. We use the publicly available LLaVA-1.5 llava15 code. Other than this modification, all training strategies follow the original code, including batch size, learning rate, and updating both the language model and the connector during fine-tuning while freezing the vision encoder. More details are available in Section I.1 and our source code [^1] https://github.com/junha1125/Lightweight-Captioner . All our experiments are conducted using two NVIDIA A6000 GPUs.
Datasets and Evaluations. The multimodal connector is first pretrained on the Caption Concept-balanced 558K llava15 , after which the model is fine-tuned on task-specific datasets such as COCO mscoco , DCI dci , or ShareGPT4V chen2024sharegpt4v , as introduced later. We consider traditional metrics such as BLEU papineni2002bleu , CIDEr vedantam2015cider , and BERTScore bertscore , as well as MLLM-as-a-Judge mllmjudge ; clair with GPT-4o-mini.
3.1 Results on Single Sentence Captioning
MS COCO Captions mscoco , the most widely used captioning benchmark, consists of 113k images and 5 captions per image. One caption typically contains a single sentence of about 10 words. The results in Table 2 show that our model outperforms previous small captioning works huang2023tag2text ; wan2024locca that utilize fewer than 1B parameters. Specifically, our model achieves CIDEr scores that are 6.9 higher than SmallCap ramos2023smallcap , which also uses OPT-125M as its LM backbone. Considering that our model does not involve any newly proposed methods, these results show surprisingly strong performance. We further discuss these results in Section E.1. Furthermore, we compare our model against MLLM generalists chen2023pali ; bai2023qwenvl . Despite requiring significantly fewer computational resources, our model achieves comparable performance to these larger counterparts.
MS COCO mscoco
model venue # data # params B@4 papineni2002bleu MET meteor CIDEr vedantam2015cider BERT bertscore CLAIR clair GPT mllmjudge
G InstructBLIP dai2023instructblip NeurIPS23 130M> 8.2B 38.0 29.4 127.8 69.1 - -
Unified-IOXL lu2023unifiedio ICLR23 130M 7.3B 37.0 29.5 123.6 68.2 - -
Shikra chen2023shikra arXiv23 - 7.2B - - 117.5 - - -
[…]
ViPCap vipcap AAAI25 4M 225M 37.7 28.6 122.9 - - -
[…]
ShareGPT4V chen2024sharegpt4v & DCI dci
model venue # data # params B@4 papineni2002bleu CIDEr vedantam2015cider BERT bertscore CAPT dong2024benchmarking CLAIR clair GPT mllmjudge
[…]
LLaVA-1.5 llava15 CVPR24 1M 7.3B 10.6 36.1 36.6 48.6 $\pm$ $\pm$
EyesWideShut eyeswideshut CVPR24 1M 7.6B 11.6 36.5 37.2 49.0 - -
[…]
SmallCap* ramos2023smallcap CVPR23 0.5M 450M 14.5 20.1 28.9 23.3 - -
[…]
GLaMM rasheed2024glamm
model venue # data # params B@4 papineni2002bleu CIDEr vedantam2015cider BERT bertscore CAPT dong2024benchmarking CLAIR clair GPT mllmjudge
[…]
Table 2: Comparison of captioning performance. We evaluate our specialist and existing captioning models. Despite not introducing any newly proposed methods, our model achieves strong performance. Here, ‘G’ represents a MLLM generalist, while ‘S’ denotes the small captioning model. ‘*’ indicates models that are fine-tuned in this study, as they were not trained for detailed captioning tasks.
Figure 3: Qualitative results on detailed captioning. Despite its small size, our specialist generates effective image captions, with additional examples shown in Figure 15. For comparison, results from generalist models are provided in Figure 12 and Figure 13.
3.2 Results on Detailed Captioning
For this experiment, we fine-tune our model on a combination of the DCI dci and ShareGPT4V chen2024sharegpt4v datasets, and additionally utilize the GLaMM rasheed2024glamm dataset. We then evaluate the model on the test splits of these datasets. Further details are provided in Section G.2.
We initially anticipated that our specialist would struggle with this challenging task, given that OPT-125M possesses limited language processing capabilities, presumably due to its small parameter count, as shown in Table 4. However, the results in Table 2 and Figure 3 reveal that our model performs surprisingly well on detailed captioning tasks, contrary to our expectations. It should be noted that the vision encoder remains frozen while only OPT-125M is updated, meaning that detailed caption generation is largely handled by OPT-125M, with the vision encoder offering visual cues.
Figure 4: Generation examples from OPT-125M. Surprisingly, despite its limited capacity, our lightweight specialist performs image captioning effectively.
Results Summary.
Although MLLMs rely on the complex reasoning abilities of LLMs for tasks such as VQA and instruction following, our experiments indicate that captioning—primarily involving the enumeration of factual objects—does not significantly require such capabilities. This observation suggests that a lightweight captioning model can serve as an efficient and practical alternative for captioning-based applications.
Note 1: Factual image captioning does not significantly require the complex reasoning abilities of LLMs and can be handled by a lightweight language model.
Figure 5: Visualization of attention maps for highlighted words. The left figure shows attention during single-pass caption generation, while the right figure illustrates the refinement process with one more glance guided by the initial caption.
Figure 6: Image reconstruction using visual representations from CLIP CLIP . Results are shown for last-layer features (i.e., LF) and multi-layer features (i.e., MF), where MF corresponds to representations extracted from layers 13, 18, and 23 of the 24-layer OpenAI CLIP ViT-L/14@336.
Despite its impressive performance, our model shares similar limitations with other MLLMs cambrian ; lu2024deepseekvl , occasionally producing incorrect predictions. Prior works have interpreted these errors from two perspectives: (i) the hallucination hypothesis, where LLMs, during response generation, describe non-existent elements in the image bai2024hallucination ; li2023evaluatingobjecthallucinationlarge ; (ii) the visual blindness hypothesis, where the visual encoder generates visual representations which are unclear or ambiguous eyeswideshut . Given that our model primarily consists of parameters dedicated to the vision encoder (e.g., 67% of total parameters), we hypothesize that the observed errors are more likely a result of visual blindness. In the following section, we provide an in-depth analysis of this issue and introduce our proposed solution.
4 Why Sharp-Eyed Refinement?
In this section, we conduct a deeper investigation into our captioning specialist, formulating two hypotheses and validating them through toy experiments and visualizations. We then present our captioning framework to address these issues.
Analysis 1.(Hypothesis.) We propose a first hypothesis: they fail to focus on critical regions (e.g., entities, attributes, relations), instead maintaining a blurry view of the entire image. One possible cause lies in the current captioning framework, which forces the model to process the entire image and produce a caption in a single pass. This is analogous to asking a human to describe an image in one glance, without attending to specific details. (Setup.) To investigate this issue, we analyze where in the image the model attends during the generation of specific words. The attention maps are visualized using code adapted from API yu2024attention , with further implementation details provided in Section I.3. (Result.) As illustrated in Figure 5, the model often exhibits diffuse attention and fails to focus on specific regions. These observations motivate the development of approaches that guide the captioning process toward more effective attention mechanisms.
Note 2: Single-pass captioning can lead models to adopt broad attention, neglecting critical visual regions necessary for precise description.
Analysis 2.(Hypothesis.) Eyes Wide Shut eyeswideshut demonstrates that MLLMs suffer from visual deficiencies, as vision encoders such as CLIP CLIP provide ambiguous representations that weaken performance. Based on this, we hypothesize that the widely used CLIP features may not provide sufficiently informative representations for generating accurate captions. (Setup.) Unlike prior studies, we examine this issue by reconstructing images from their visual features. The quality of the reconstructed images indicates how much visual detail the features preserve and, consequently, their effectiveness in supporting captioning tasks. For this experiment, we directly employ the decoder architecture from Masked Autoencoder (MAE) mae . We keep the CLIP encoder frozen while training only the MAE decoder on the MS-COCO images. (Result.) As illustrated in Figure 6, the images reconstructed from CLIP embeddings exhibit significant deviations from the original images and appear less clear. This result indicates that such visual representations may not provide sufficiently detailed visual information for captioning.
Note 3: CLIP’s final-layer features provide only coarse-grained visual information, limiting their suitability for fine-grained captioning tasks.
Our proposed solution.Building on the above insights, we propose a novel captioning framework, Sharp-Eyed Refinement (SeR), illustrated in Figure 1 and Figure 7. The framework is inspired by the human description process, where one first takes an initial glance and then sharpens the eyes to capture finer details. Concretely, the model generates an initial caption that is subsequently refined by the LM through our DeepLens. This refinement process resembles approaches in LLM research such as self-refinement selfrefine ; le2022coderl ; pan2023automatically , where models improve their outputs by revisiting earlier generations. By extending this idea to the multimodal domain, we uniquely integrate visual features into the refinement step, and, to the best of our knowledge, this constitutes the first work in this direction. Further implementation details are provided in the following section.
5 Inside Our Framework: DeepLens and Learning Strategy
Figure 7: Inference and training pipeline for SeR. The model is trained in two fine-tuning stages: (1) initial caption generation and (2) caption refinement. At inference, both an initial caption and a refined caption are produced, as shown by the solid and dashed lines in the rightmost figure.
In this section, we take a closer look at our framework and its effectiveness. At the core lies DeepLens, a multimodal connector that, unlike prior designs llava ; li2023blip2 ; bai2023qwenvl , leverages two new inputs: (i) previously generated captions and (ii) multi-layer features from the vision encoder. These inputs guide our model to attend to salient regions, enrich visual representations, and enable effective caption refinement.
Look at what matters.Imagine receiving an image along with the caption “A cat relaxing on a brown chair” and being instructed to refine it. In such a scenario, humans naturally focus on key elements like the cat and the chair. We attempt to employ a similar mechanism by providing the initially generated caption as an input for DeepLens. This design enables the module to identify and attend to the relevant visual regions. Similar to the experiment in Section 4, we also inspect the attention maps while the model takes a second glance using the previous captions, as shown in Figure 5. We observe that the model focuses more precisely on regions corresponding to specific words, indicating that it attempts to look at what matters in the image.
Look in detail.To mitigate the limitation of coarse-grained visual features, some studies cambrian ; eyeswideshut ; fromcliptodino proposed using additional vision encoders, but this comes at the cost of additional model parameters. For example, Interleaved MoF eyeswideshut adds 300M parameters for DINOv2 oquab2024dinov —a substantial 60% increase relative to our 500M-parameter model. Rather than introducing additional encoders, we focus on maximizing the utility of the existing vision encoder. To this end, we design DeepLens to leverage multi-layer features, expecting them to provide semantically richer representations that enable more accurate captioning, inspired by FPN fpn . As demonstrated in the reconstruction experiment (Figure 6), utilizing these features yields clearer images. This suggests that the multi-level visual features indeed contain fine-grained information, thereby helping the LM recognize visual content in greater detail.
Note 4: Encouraging attention to key regions via the initial caption, along with multi-level utilization of a single vision encoder, can be an effective and efficient strategy.
Figure 8: Examples of captions during SeR. Initial outputs occasionally include inaccuracies, which are corrected through the refinement process. More results can be found in Section F.
Fine-tuning strategy.Assume we have a training set $X=\{x_{1},\dots,x_{N}\}$, where each $x_{k}=(i_{k},c_{k})$ consists of an image $i_{k}$ and its corresponding ground-truth caption $c_{k}$; $N$ denotes the total number of such pairs. As illustrated in Figure 7, we adopt a two-stage fine-tuning strategy. In the first stage, we focus on enabling the model to generate an initial caption by following the training procedure of LLaVA. Specifically, we perform supervised fine-tuning on the MLP connector and the lightweight language model using this dataset.
In the second stage, we focus on teaching the model the refinement process. A naïve approach would be to directly use the momentary generated captions as initial inputs and the true captions $c_{k}$ as targets for refinement, but this can be misleading. Suppose our captioning specialist were to initially produce something like ‘a brown table in front of a window’ while the true caption is ‘a cat sitting on a chair’; these two descriptions share little in common and provide no meaningful signal for progressive refinement. We believe that training on such disparate pairs would encourage the model to ignore the provided initial caption and generate a wholly new one, rather than learn to refine it. We provide additional explanations in Section E.5. To address this challenge, we construct pseudo‐initial captions $\hat{c}_{k}$ that remain close to their corresponding ground-truth $c_{k}$. In practice, we leverage GPT-4o-mini achiam2023gpt to slightly alter entities, attributes, or relations within $c_{k}$. For example, given the true caption ‘a cat sitting on a chair’, the pseudo-initial caption might become ‘a dog sitting on a chair,’ preserving the overall structure while introducing minor variations.
Leveraging these pseudo-initial captions, each pair $(i_{k},\hat{c}_{k})$ is fed into DeepLens together with the language model, and optimization is performed using a next-token prediction loss (i.e., supervised fine-tuning). Our fine-tuning schedule consists of 10 epochs for the first stage and 2 epochs for the second stage. Further details of the training setup and hyperparameters are provided in Section I.1, and we describe the generation strategy for $\hat{c}_{k}$ in Section G.1, along with an ablation study on refinement data in Section E.5.
Rationale for refinement fine‑tuning.Let $\pi_{\theta}$ denote the language model. Each pseudo‑initial caption $\hat{c}_{k}$ deviates from the ground truth $c_{k}$ at only a few token positions $E_{k}=\{t\mid\hat{c}_{k,t}\neq c_{k,t}\}$. Under the sequence‑level cross‑entropy objective
$\mathcal{L}(\theta)\;=\;-\!\mathbb{E}\Bigl[\sum_{j}\log\pi_{\theta}\!\bigl(c_{k,j}\mid i_{k},\hat{c}_{k},c_{k,<j}\bigr)\Bigr],$ (1)
gradients are likely to be primarily concentrated on tokens in $E_{k}$, leading to a form of targeted optimization, in which the model retains the correct parts of $\hat{c}_{k}$ while rewriting only the erroneous ones. Writing $\Delta_{k}(\theta)=\log\pi_{\theta}(c_{k}\mid i_{k},\hat{c}_{k})-\log\pi_{\theta}(\hat{c}_{k}\mid i_{k},\hat{c}_{k}),$ we have $\mathcal{L}(\theta)\propto-\mathbb{E}[\Delta_{k}]$; hence, minimizing $\mathcal{L}$ is equivalent to maximizing the expected margin $\Delta_{k}$, thereby directly increasing the likelihood of the refined caption relative to its flawed precursor. In effect, each gradient step encourages DeepLens to focus on visual regions likely responsible for errors in the initial caption and guides the language model to better interpret these refined features, resulting in more accurate captions. We note that this targeted optimization resembles the philosophy of Direct Preference Optimization (DPO) dpo , if we consider $c_{k}$ and $\hat{c}_{k}$ as preferred and less-preferred responses. In contrast to DPO, which treats both responses symmetrically during optimization, our method offers a new perspective by assigning them distinct roles as input and target.
Inference strategy.Given an image $i_{k}$ and instruction $p$, the model generates an initial caption $o_{initial}$, storing intermediate ViT vit features in a buffer $B$ to avoid redundant computation. $o_{initial}$ is then fed into DeepLens and the LM to produce the refined output $o_{refined}$. Unless otherwise specified, we employ a single-step refinement. A discussion and experimental results for a repeated process are presented in Section E.6 and Section C.6, respectively.
MS COCO mscoco
model #params B@4 papineni2002bleu gain CIDEr vedantam2015cider gain CLAIR clair gain GPT mllmjudge gain
[…]
+ SeR with 1 500M 39.6 +0.2 131.9 +2.3 - - - -
+ SeR with 2 500M 39.6 +0.2 132.3 +2.7 - - - -
Single glance with 2 450M 39.7 +0.3 130.6 +1.0 - - - -
ShareGPT4V chen2024sharegpt4v & DCI dci
model #params CIDEr vedantam2015cider gain CAPT dong2024benchmarking gain CLAIR clair gain GPT mllmjudge gain
[…]
GLaMM rasheed2024glamm
model #params CIDEr vedantam2015cider gain CAPT dong2024benchmarking gain CLAIR clair gain GPT mllmjudge gain
[…]
Table 3: Quantitative results of Sharp-Eyed Refinement (SeR). The results demonstrate the effectiveness of SeR in improving caption quality. We also present an ablation study on the key inputs to DeepLens: ① initial captions and ② multi-layer features.
6 Empirical Validation
model time
LLaVA 274.49 s
[…]
Figure 9: Inference time required to generate captions for 100 streaming images.
Ablation Study of SeR. The results in Table 3 and Figure 8 demonstrate the effectiveness of SeR in improving caption quality. We further conduct an ablation study on the key inputs to DeepLens—① initial captions and ② multi-layer visual features—to examine their individual contributions. The refinement process improves initial captions by +3.9 and +2.5 CIDEr points for single-sentence and detailed captioning tasks, respectively. Although integrating only multi-layer features (i.e., using ②) yields modest gains, we observe that its effectiveness may be constrained by the single-pass generation mechanism. Moreover, our method introduces approximately 50M additional parameters for DeepLens and one extra inference step with OPT-125M. Nevertheless, as shown in Table 9, the computational overhead remains minimal relative to the substantial improvements achieved. We further validate this efficiency on a resource-constrained device, Jetson Nano, with results reported in Section C.3.
Evaluation on Long Range Video Question Answering.
Since captioning models serve as foundational components in downstream applications as mentioned in Section 2, it is important to verify their utility in a practical task. To this end, we evaluate our model on the Long-Range VQA task introduced in zhang2023simple . Following the baseline setup of Zhang et al. zhang2023simple , each captioner generates captions for multiple frames from videos spanning several minutes to hours. These captions are then fed into an LLM to answer video-related questions. As shown in Table 4, our model achieves competitive accuracy with significantly fewer parameters and faster inference than MLLM generalists, demonstrating its potential for real-world applications. The reported time includes both the captioning and LLM inference steps, assuming a 10-minute video processed at the frame level. The small specialists ramos2023smallcap ; huang2023tag2text , including ours, use weights fine-tuned on ShareGPT chen2024sharegpt4v  & DCI dci .
LLM Captioner LLaVA-1.0 llava BLIP-2 li2023blip2 LLavA-1.5 llava15 SmallCap* ramos2023smallcap Tag2text* huang2023tag2text Our specialist + SeR
Qwen2.5 14B yang2024qwen2 #params 7.3B 7.4B 7.3B 450M 900M 450M 500M
accuracy 47.6 50.6 51.1 41.8 47.1 49.3 50.8
time 29m 20s 29m 44s 29m 20s 4m 45s 7m 14s 4m 53s 5m 10s
Table 4: Evaluation on Long-Range VideoQA. We follow the baseline setup of LVQA zhang2023simple , replacing the captioner with our specialist or MLLM generalists. Captions from all frames are aggregated and fed into Qwen2.5-14B to perform the final video question answering. Higher-quality captions from our specialist contribute to competitive accuracy with significantly fewer parameters.
Note 5: Although lightweight captioners play an important role, they have received limited attention. Our proposed baseline and framework highlight their promise as efficient solutions for on-device applications.
7 Related Work
Multimodal Large Language Models (MLLMs)
have attracted considerable research attention due to their versatile applications, such as chat-bots wu2023multimodal . Early approaches integrated contrastive image-language pretrained models wang2022git with powerful LLMs, enabling complex reasoning. The development of instruction-based datasets llava and innovative training strategies li2022mplug ; ye2024mplug have further accelerated progress, substantially improving MLLM performance and broadening their capabilities. Despite these achievements, most MLLMs heavily depend on large-parameter LLMs, making deployment on memory-constrained devices unfeasible. This limitation will likely restrict access for a significant portion of users worldwide.
Image Captioning Models.
Recent advances in image captioning have improved training efficiency and descriptive fidelity. Approaches such as CaMEL barraco2022camel and SmallCap ramos2023smallcap emphasize minimizing trainable parameters by leveraging mean-teacher distillation and employing retrieval augmentation, while Tag2Text huang2023tag2text and LoCCa wan2024locca introduce novel mechanisms such as dedicated tagging and location-aware refinement to improve caption quality. Unlike existing approaches that focus on reducing trainable parameters, or rely on single-pass inference—potentially missing crucial details—our method prioritizes inference efficiency considering on-device operation and systematically addressing the limitations of single-pass generation.
Self-Refinement in LLMs.
Humans often refine their writing through iterative review to enhance clarity and precision selfrefine ; dou2024stepcoder . Recent research has applied this refinement concept to LLMs, introducing techniques. For instance, Self-Refine le2022coderl ; ranaldi2024self ; pan2023automatically enabled models to autonomously critique and iteratively enhance their outputs. Unlike such approaches limited to the LLM domain, our method introduces refinement in a multimodal context, guided by both language and vision.
8 Conclusion & Broader Impact
[…]
Broader Impact.
We demonstrate the potential of small-scale specialists, highlighting their surprising effectiveness. As a next step, future research may explore iterative refinement strategies, integrating external tools (e.g., zoom or crop, as in GPT-4o openai2024o3o4 ), and designing a unified multimodal connector for both the initial and secondary glance. More research questions that aim to enhance both captioning performance and efficiency in real-world applications are provided in Section E.6.
Acknowledgments
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (RS-2022-00143911,AI Excellence Global Innovative Leader Education Program)
Appendix
Appendix A Summary of Our Work
Motivation
  1. 1.
    Image captioning serves as the cornerstone of diverse applications such as assistive devices.
  2. 2.
    However, running multimodal large language models for this task is challenging for edge devices.
  3. 3.
    We highlight the practical importance of lightweight captioning.
What We Have Demonstrated?
  1. 1.
    (Note1) Factual image captioning does not significantly require the complex reasoning abilities of LLMs and can be handled by a lightweight language model.
  2. 2.
    (Note2) Single-pass captioning can lead specialists to adopt broad attention, neglecting critical visual regions necessary.
  3. 3.
    (Note3) CLIP’s final-layer features provide only coarse-grained visual information, limiting their suitability for fine-grained captioning tasks.
Novel Caption Refinement Framework
  • What is it? Humans first take in the overall scene, then glance at specific regions to notice finer details. Our Sharp-Eyed Refinement framework mimics this human tendency, allowing the captioning specialist to revise and improve initial descriptions.
  • (Note4) Encouraging attention to key regions via the initial caption, along with multi-level utilization of a single vision encoder, can be an effective and efficient strategy.
  • Why novel? To the best of our knowledge, this is the first work to apply multimodal refinement that jointly utilizes visual features and the model’s initial textual output.
    • $\circ$
      NLP community: Humans often refine their writing, and coders revise their code through iterative review. Recent research has applied this refinement/correctness concept to LLMs [51, 60, 82, 32]. Unlike such approaches limited to the LLM domain, our method introduces refinement in a multimodal context.
    • $\circ$
      Multimodal LLMs community: Existing models [44, 3, 97] heavily rely on the complex reasoning capabilities of LLMs and typically generate outputs in a single pass. We uniquely demonstrate the effectiveness of revisiting and refining its own textual outputs.
    • $\circ$
      Visual Blindness community: Several studies [65, 66] believe the visual encoder is a critical bottleneck and adapting multiple vision encoders. However, this requires additional parameters. Since our research focuses on a lightweight model, we maximize the utility of the existing vision encoder through multi-level utilization.
    • $\circ$
      Lightweight captioning community: This field has been gradually declining since the remarkable capabilities of LLMs were discovered. In this work, we revisit the practically important yet underexplored topic of lightweight captioning models.
        Unlike ours, some methods [49, 59, 73] utilize prior captions obtained through a heavy image-text retrieval process. For example, SmallCap performs similarity matching between a given image and 500,000 candidate captions at each iteration.
        Unlike ours, several works [25, 48] incorporate object detection or tagging procedures. However, despite these extra components, their captioning performance has been limited.
        Unlike ours, which refines text output generated via a full forward pass, certain works [49, 12] adopt Diffusion Transformers and denoise latent text embeddings. Moreover, they provide little motivation or analysis as to why such a process is important from the perspective of utilizing visual cues more effectively.
Broad Comparisons
  1. 1.
    This is the first paper to demonstrate the amazing capability of a lightweight captioning model beyond single-sentence captioning, including not only detailed captioning but also long-form VideoQA.
  2. 2.
    We compared not only existing small captioning models but also large multimodal models, demonstrating the potential of our model as a strong visual specialist.
  3. 3.
    Through these extensive experiments, (Note5) we show that our proposed baseline and framework have the potential of lightweight captioning for on-device applications.
Positive Influences
  1. 1.
    Industry: Our model can be easily applied in industrial scenarios such as assistive devices for the visually impaired or navigation robots. Our model runs successfully on actual edge devices like Jetson Nano.
  2. 2.
    Academy: For researchers with limited access to large-scale GPUs who are unable to operate LLMs, our factual image captioning model offers a promising baseline.
Appendix B Additional Related Work
B.1 Visual Blindness in VLMs.
Despite significant advancements, MLLMs still face limitations in their visual capabilities, hindering their practical applications. Eyes Wide Shut [66] demonstrated that even GPT-4V [1] struggles with basic visual questions. Research on this topic typically points to two main sources of failure: one relates to the language decoder, which can hallucinate details not present in the image [5], while the other focuses on the visual encoder, which may provide ambiguous visual information. Several studies, including Cambrian [65], believe the visual encoder is a critical bottleneck. We also concentrate on the visual issue, particularly within the context of our lightweight model, where the visual encoder accounts for a significant portion of the parameters. Furthermore, we introduce a novel operational framework to improve visual grounding.
Appendix C Additional Experiments
C.1 When GPT meets ‘one more glance’
In this work, we suggest that allowing the model to take one more glance, guided by the initial caption and aligned visual features, offers an effective approach, as discussed in Note 4. We further explore how OpenAI’s GPT models, which have demonstrated remarkable performance, might be influenced by our refinement framework. While our framework consists of two components, taking one more glance and refining with sharp eyes through DeepLens, we apply only the former as modifying their internal architecture is not feasible. The results are shown in Figure 10, where GPT-4-turbo is used. As a commercial model with a sufficiently large number of parameters, it often produces an accurate initial caption without noticeable errors, but we still observed occasional inaccuracies in its outputs. To explore whether refinement could help, we instructed the model to take one more glance and found that GPT can generate a refined caption by re-examining the given image. OpenAI’s GPT-o3 reflects an awareness of this idea through the introduction of the Thinking with Images [53]. However, this concept remains less explored in the research community, suggesting the need for further investigation.
C.2 When larger specialists meet ‘one more glance with sharp eyes’
Building on our lightweight specialist, we naturally ask: since our framework and module offer a scalable methodology, can they also be effective when applied to larger multimodal models? To investigate this, we scaled the language model from OPT-125M to OPT-1.3B and also utilized LLaMA-2-7B. Specifically, we fine-tuned these models on the ShareGPT and DCI training datasets to serve as captioning specialists (i.e., single pass). Next, we incorporated our Sharp-Eyed Refinement (i.e., SeR) framework with DeepLens and further fine-tuned them using the datasets described in Section 5. As shown in Table 5, our framework consistently improves caption accuracy across different models. In particular, we observe gains of 1.2 and 0.9 points in CAPT [18] evaluation for OPT-1.3B and LLaMA-2-7B, respectively. Additionally, the results indicate that using larger language models to build captioning specialists leads to improved performance, observed in prior LLM scaling studies [30, 79, 24]. Despite the gains observed with larger models such as OPT-1.3B and LLaMA-2-7B, we note that their considerable size (roughly 10x and 56x larger than OPT-125M, respectively) may hinder real-world applicability. It is therefore crucial to continue exploring captioning systems that balance effectiveness with efficiency.
Figure 10: When GPT meets ‘one more glance’. Example outputs from OpenAI’s GPT before and after being prompted to take one more glance. The refined captions illustrate its potential responsiveness to visual re-examination.
ShareGPT4V [11] & DCI [70]
vision total CIDEr [71] CAPT [18]
Language model encoder LoRA [23] # params single pass + SeR single pass + SeR
OPT-1.3B [93] CLIP ViT-L $\times$ 1.6B 50.2 53.1 (+2.9) 49.0 50.2 (+1.2)
LLaMA-2-7B [67] CLIP ViT-L 7.3B 57.3 61.7 (+4.4) 52.7 53.6 (+0.9)
Table 5: Captioning performance of larger specialists with and without Sharp-Eyed Refinement (SeR). SeR consistently improves accuracy across model scales, including OPT-1.3B and LLaMA-2-7B.
C.3 Evaluation on an Actual Edge Device
To highlight the practical value of our approach, we present an evaluation of our lightweight captioner on edge devices, where large-scale MLLMs are typically infeasible to deploy. Our core motivation is rooted in the observation that while MLLMs are powerful, they are computationally intensive and difficult to run in resource-constrained environments. In contrast, lightweight captioners remain underexplored despite their potential for real-world applications.
Here, we demonstrate that lightweight captioners can be effectively deployed on edge devices such as RTX 3090 and Jetson Nano. We evaluate captioning performance on MS COCO and ShareGPT4V&DCI datasets, and additionally measure inference time, memory usage, and power consumption. All models were executed for 100 iterations with a batch size of 1. The results in Table 6 show that our method consistently performs well across devices, and notably, remains fully operational on setups where models like LLaVA-1.5 fail to run. These findings support the deployability of our framework in edge scenarios and encourage further exploration of lightweight captioning for real-world assistive technologies.
C.4 Comparison on different learning strategies
As writers frequently revise their drafts to improve clarity or tone, and programmers refactor code to enhance structure, we demonstrate that allowing multimodal models to refine their own outputs through our framework can be an effective strategy. In this subsection, we examine how our lightweight specialist performs under different learning strategies. The first strategy involves training with as much data as possible. Here, we train the model on a combined set of COCO, ShareGPT, DCI, and GLaMM to maximize data exposure (i.e., more data). The second strategy explores distillation using teacher-generated captions. Specifically, we train our model on captions produced by LLaVA-1.6-34B [45] (i.e., distillation). Our original strategy was straightforwardly fine-tuning the model exclusively on task-specific data (i.e., origin). The results are summarized in Table 7. For the ‘more data’ strategy, we observe limited performance, likely due to reduced task alignment caused by mixing heterogeneous datasets. In contrast, the ‘distillation’ strategy demonstrates that learning from a strong teacher can improve performance even in a single-pass setting. Notably, when combined with our refinement framework (SeR), it yields further gains, suggesting that distillation and SeR can be effectively integrated. Additionally, as noted in Section 5, the datasets we used share similarities with those in Direct Preference Optimization (DPO). Investigating the impact of training our model using reinforcement learning methods such as DPO could be a promising direction for future work.
Resource (RAM) Data Model Mem. Inf. time Power. B@4 [56] CIDEr [71] CLAIR [8] GPT [9]
Jetson Nano (4G) All LLaVA-1.5-7B out-of-memory N/A N/A N/A N/A N/A N/A
[…]
Table 6: Evaluation results across different hardware resources and datasets.
ShareGPT4V [11] & DCI [70]
Metric origin +SeR gain more data distillation +SeR gain
CIDEr [71] 40.5 43.6 +3.1 36.8 42.6 43.6 +2.0
CAPT [18] 45.9 48.4 +2.5 43.2 46.5 47.4 +0.9
Table 7: Performance comparison of our lightweight captioner under different learning strategies: origin, more data, and distillation. Distillation from a strong teacher, especially when combined with Sharp-Eyed Refinement (SeR), leads to the best results.
C.5 Efficacy of Multi-Level Features with Other Vision Encoders
We investigate whether our strategy of selecting multi-level features can generalize to vision encoders beyond CLIP [57], which is our original setup. As shown in Table 8, the approach consistently improves performance across multiple encoders. Specifically, both SigLIPv2 [68] and DINOv2 [54] exhibit performance gains when equipped with multi-level features, demonstrating that our method is not limited to CLIP-based models. Interestingly, CLIP with multi-level features outperforms the combination of CLIP and DINOv2 in several evaluation metrics, while maintaining greater parameter efficiency. On the other hand, DINOv2 alone yields relatively lower performance, likely due to weaker alignment with language features. For these experiments, we pair OPT-125M with each vision encoder and fine-tune the resulting models on the ShareGPT [11]&DCI [70] datasets. Subsequently, we apply our Sharp-Eyed Refinement framework using the selected multi-level features and perform second fine-tuning with the datasets described in Section 5 of the main paper.
Vision encoder #params indices of selected layers CIDEr [71] CLAIR [8] GPT [9]
CLIP [57] 300M {23} 42.8 55.8 2.78
300M {13, 18, 23} 43.3 57.7 3.02
CLIP [57]+DINOv2 [54] 600M {23} + {23} 42.9 57.3 3.02
SigLIPv2 [68] 300M {23} 43.0 56.2 2.88
300M {15, 23} 45.9 57.7 3.05
300M {13, 18, 23} 45.5 58.2 3.07
DINOv2 [54] 300M {23} 32.8 48.6 2.55
300M {13, 18, 23} 33.0 50.1 2.66
Table 8: For the captioning specialist, we evaluate different vision encoders and the corresponding selected layer indices.
C.6 Iterative Sharp-Eyed Refinement
To explore the potential of iterative refinement beyond a single pass, we conduct additional experiments using captioning specialists with larger model capacity. This decision is motivated by our initial hypothesis that the constrained capacity of small models may have hindered the effectiveness of multi-step refinement. As shown in Table 9, applying multiple refinement iterations to the smallest language model, OPT-125M, results in only limited improvements. In contrast, we observe that multi-step refinement leads to performance gains when used with higher-capacity captioners. These findings underscore the potential of scaling up refinement strategies, as similarly observed in recent LLM-based studies [51, 60, 82, 32]. We plan to further investigate this direction in future work.
For implementation, we construct larger captioning specialists based on two configurations: LLaMA-2-7B + CLIP-ViT-Large and OPT-1.3B + CLIP-ViT-Large. These models are initially fine-tuned on the ShareGPT and DCI training datasets via single-pass training. We then apply the Sharp-Eyed Refinement (SeR) framework with DeepLens in multiple iterations, followed by additional fine-tuning using the datasets described in Section 5.
Language model vision encoder Benchmarks Eval. matrix initial cap. refinement refinement *2 refinement *3 refinement *4
OPT-125M CLIP ViT-L MS COCO CIDEr [71] 129.6 133.5 133.8 133.1 133.4
OPT-1.3B 133.7 135.7 138.1 138.1 138.5
LLaMA-2-7B 137.4 141.4 141.6 141.9 141.7
[…]
Table 9: Multiple iterative refinements with larger language models. This experiment demonstrates better refinement capability as model capacity increases.
Appendix D DeepLens
D.1 Architectural details
We have shown that our refinement framework can improve caption quality. At the core of it lies DeepLens, which we hypothesize plays a critical role by aligning multi-level visual features with semantic cues from the initial caption. As illustrated in Figure 11, DeepLens takes two new inputs: ① the previously generated caption, represented as token embeddings, and ② multi-layer features extracted from the vision encoder. Let $N$, $d$, and $m$ denote the number of visual tokens, the embedding dimension, and the number of selected layers, respectively. The selected features are concatenated along the channel dimension to retain hierarchical visual information. These fused inputs are then passed through Transformer layers following the BERT architecture [17]. The resulting output is provided to the language model, enabling refinement conditioned on a focused and semantically grounded visual representation.
Figure 11: (i) Illustration of the DeepLens architecture. This is designed for integrating ① token embeddings from the initial caption and ② multi-layer visual features for refinement. (ii) and (iii) Ablation study of multimodal connector designs. Each configuration corresponds to evaluation in Table 10.
D.2 Ablation study
We conduct a series of ablation experiments to analyze the design choices behind DeepLens, as summarized in Table 10. The corresponding model variants are illustrated in Figure 11. Our study spans three key aspects: connector selection, refinement configurations, and the layer indexing strategy for visual feature extraction. (I) We begin by comparing different multimodal connectors used in the initial caption generation phase. While prior works have explored a range of designs, such as Cross-Attention [59, 6], Q-Former [39], and Transformer-style modules, we find that the simple MLP connector from LLaVA [43] offers competitive performance. Notably, (b) and (f), which adopt multi-level features and a BERT-style Transformer block respectively, perform slightly better than the MLP-based connector. However, given the marginal gains, we retain the lightweight MLP structure for efficiency. (II) Next, we evaluate several connectors for the refinement phase (SeR). Among them, combining our base configuration (a) with structure (A), which utilizes both proposed inputs, yields strong performance. Although we also applied (A) to configuration (b), the improvement was marginal, so we did not adopt it as our baseline. (III) Finally, we investigate the impact of selecting different sets of layers from the ViT encoder. We test multiple combinations and find that using a diverse set of layers {13, 18, 23} results in the highest performance.
(I) connectors for initial generation
(a) (b) (c) (d) (e) (f)
129.6 ✓ 130.6 125.9 127.7 122.9 130.9
(II) (a) + connectors for SeR (b) + con.
(A) (B) (C) (D) (A)
133.5 ✓ 131.9 132.1 131.9 133.6
(III) indexes of selected layers in ViT
{23} {13, 23} {15, 23} {15, 19, 23} {13, 18, 23}
131.9 133.0 133.0 133.3 133.5 ✓
Table 10: Ablation results evaluating connector types, refinement configurations, and ViT layer selections in DeepLens. Architectural variants are illustrated in Figure 11 (b) and (c).
Appendix E Discussions
E.1 Why do previous small models fall short?
The development of captioning models can be divided into two eras: before and after the adoption of LLMs. Prior to the integration of LLMs, many approaches relied on architectures that did not involve a significant number of parameters. For example, CaMEL [6] and SmallCap [59] employed GPT-2 models with 125M to 350M parameters, while Tag2Text [25] and LoCCa [72] used BERT models ranging from 300M to 900M parameters. The emergence of models such as ClipCap [52], BLIP [38], and LLaVA [44] shifted the direction of the captioning field. With growing interest in the impressive capabilities of LLMs, much of the research has focused on building MLLMs.
In contrast to this trend, we analyze small captioning models and uncover the following: Aforementioned small captioning models typically adopt an architecture where visual features are injected via cross-attention. Unfortunately, when using a small language model, this design does not appear to be particularly beneficial. Empirically, this limitation is reflected in the performance of structure (c) in Figure 11, which yields a relatively low score of 125.9 as shown in Table 10. Had the field not been so heavily influenced by the rise of LLMs, such limitations in small model architectures might have been identified and addressed earlier.
We adopt the LLaVA architecture and inject visual features into the self-attention inputs, as shown in Figure 11 (a). We show that this simple setup surprisingly delivers stronger performance (i.e., Note 1 in Section 3) and contributes to demonstrating that small models can be viable for practical applications. We hope this encourages a shift toward reducing dependence on LLMs in specific tasks like captioning, and sparks further interest in the development of lightweight yet effective models.
E.2 Limitations of Existing Multimodal Large Language Models
We examine the broad challenges faced by existing multimodal large language models (MLLMs), particularly in addressing the issue of visual blindness. Prior works such as Eyes Wide Shut [66] and Cambrian [65] have identified and attempted to mitigate this issue through the use of multiple vision encoders, including DINOv2 [54], SigLIPv2 [68], and CLIP [57]. However, as illustrated in Figure 12, even models with large parameter sizes continue to exhibit difficulty in generating consistent long-form captions in complex, multi-object scenes. We also conduct experiments using two recent capable MLLMs, LLaVA-Next [37] and LLaVA-OneVision [35]. Despite incorporating the sophisticated approach which is to split an image into grids and utilize all the features from each grid, both models continue to suffer from incorrect captioning results.
These findings suggest that visual blindness is a persistent limitation across model scales and architectures. In this context, we believe that our Sharp-Eyed Refinement framework—which encourages attention to key regions via initial captions and leverages multi-level features from a single vision encoder—is a meaningful and efficient step toward addressing this challenge.
E.3 Limitations of lightweight captioners
While we have demonstrated the potential of small models in captioning tasks, the use of lightweight LMs unavoidably introduces some limitations. In particular, we observe that the model occasionally suffers from issues such as repetitive phrasing, reduced fluency, limited OCR capability, and a lack of general world knowledge. Examples of these cases are provided in Figure 13. These limitations may stem from two primary factors: (i) the small number of parameters, which can restrict the model’s capacity for complex reasoning and language generation [30], and (ii) the limited scale and quality of training data, as our model was trained on approximately 500K image-caption pairs from ShareGPT-4V [11] which contains machine-generated captions. A natural direction for future work is to investigate how far the capabilities of small models can be scaled with access to larger and higher-quality training datasets. In addition to the results in Section E.2, we observe similar issues in larger models such as LLaVA-1.5, suggesting that these challenges remain unresolved [5, 66] and require deeper investigation.
E.4 Limitations of existing evaluation methods
To ensure fair comparison across captioning models, we adopt seven evaluation metrics: BLEU@4 [56], METEOR [16], CIDEr [71], BERTScore [94], CAPTURE [18], CLAIR [8], and MLLM-as-judge [9] (i.e., GPT-based evaluation). For CLAIR and MLLM-as-judge, we randomly select 100 image samples and evaluate them using 10 different seeds to compute both mean and standard deviation. The values reported alongside the $\pm$ symbol indicate the standard deviation across seeds. Despite these efforts, we find that current evaluation methods do not always align well with human judgment. For instance, certain models perform better under specific metrics but worse under others, leading to inconsistent rankings. Moreover, small specialists including ours sometimes exhibit deficiencies in fluency, as discussed in Section E.3, which existing metrics struggle to capture. As we examined, MLLM-as-judge yields relatively reliable results. In contrast, CLAIR showed higher variance across runs, limiting its consistency.
These observations highlight the need for further research into robust captioning evaluation. In particular, future work should account for multiple dimensions of quality, including fluency, coherence, faithfulness, relevance, informativeness, and completeness. Moreover, moving beyond traditional n-gram-based matching, reliable evaluation methods powered by MLLMs (e.g., OpenAI GPT) are needed. We believe such advancements can support greater trust in captioning-based applications.
Figure 12: Despite their size, existing MLLMs struggle with visual blindness, especially in complex scenes. We hope our approach offers a meaningful step toward alleviating this issue.
Figure 13: Despite their efficiency, small models suffer from various limitations. Further research is required to assess how high-quality data can compensate for their weaknesses.
E.5 Impact and Ablation of Pseudo-Initial Captions in Stage-2 Fine-tuning
We first provide a discussion, with examples, on how the inclusion of pseudo-initial captions leads to better supervision during training. We highlight the following points: (i) The model is trained to generate the ground-truth (GT) caption given both the image and the pseudo-initial caption as input: $I+\hat{c}\;\rightarrow\;\text{Our model}\;\rightarrow\;c$. (ii) If the pseudo-initial caption is generated following the strategies in Section G.1, then it is unlikely to include too many differing tokens from the GT caption.
pseudo-initial caption $\hat{c}$ different tokens $E={t\mid\hat{c}(t)\neq c(t)}$ in Section 5
A woman in a room with two dogs two / dogs
A cat sitting on a chair in front of the window. sitting / on a chair / in front of the window
Table 11: Token-level differences between pseudo-initial and ground-truth captions.
Suppose we have “A woman in a room with a cat.” as the $c$ and $\hat{c}$ in Table 11.
In such cases, if each pair of pseudo-initial caption and GT caption contains a reasonable number of overlapping tokens, the model is encouraged to refer to the pseudo-initial caption and selectively revise the differences to produce the GT caption. Conversely, if the pair contains too many differing tokens, the model is likely to avoid relying on the pseudo-initial caption and instead generate the GT caption independently. This mechanism helps the model refine incorrect parts rather than regenerate the caption from scratch, thereby mitigating misleading supervision.
As an additional experiment, we fine-tune our captioner via Stage 2 as in Figure 7 using pseudo-initial captions generated under the following four conditions: (Data 1) Initial captions generated by the captioner fine-tuned via Stage 1, (Data 2) Pseudo-initial captions that are significantly different from the GT caption, (Data 3) Pseudo-initial captions with minor modifications from the GT caption, which is our original strategy, and (Data 4) Pseudo-initial captions consisting of (i) one generated as in Data3, and (ii) the GT caption itself (e.g., GT: “A woman in a room with a cat”; pseudo-initials: “A boy in a room with a dog” and “A woman in a room with a cat”).
Detailed captioning CIDEr [71] gain CAPT [18] gain GPT [9] gain
[…]
Table 12: Detailed captioning results on ShareGPT4V [11] & DCI [70] with fine-tuning on four types of pseudo-initial captions. The comparison highlights how different levels of overlap or inclusion of GT captions affect model performance.
Table 12 demonstrates that Data 1 provides reasonable gains within our framework, while Data 2 leads to the anticipated issue where the model tends to ignore the initial caption and generate a new one. Interestingly, Data 4 performs comparably to Data 3, and together with the results from Data 1, indicates that our framework is robust to moderate variation in pseudo-initial caption quality.
E.6 Further Research Questions
To guide future exploration, we outline several research directions that may advance the field of lightweight captioning and multimodal learning more broadly:
  1. 1.
    What is the minimal model size required for a captioning specialist to be practically useful in real-world assistive technologies? Where does the performance-efficiency trade-off stabilize?
  2. 2.
    Although supervised fine-tuning is commonly used in multimodal training, recent work has shown the benefits of reinforcement learning methods in the LLM field [67]. Can similar reward-based training be applied to enhance captioning models?
  3. 3.
    The LLaVA framework remains a popular baseline due to its simplicity and open-source accessibility. However, could a unified architecture, such as a fully integrated Transformer-based multimodal model, serve as a better foundation for building practical captioners?
  4. 4.
    Multimodal models heavily rely on the quality of the representations produced by vision encoders. To what extent can we trust visual features from models like SigLIP [90] or MAE [92], especially for fine-grained captioning tasks?
  5. 5.
    Similar to self-refinement in LLM research [51], how can we enable our framework to iteratively improve its own outputs?
  6. 6.
    Can our framework be effective in simple VQA settings? Is an LLM necessary in such setups?
  7. 7.
    Do potential biases (e.g., gender, occupation, and age) exist in captioning models, and if so, how can we address them together with cultural contexts and deployment domains such as indoor, outdoor, and robotic environments? Could domain adaptation with user feedback serve as a solution?
We consider these questions promising directions for future research.
Figure 14: Additional results corresponding to the analyses in Figure 5 and Figure 6. It illustrates diffuse attention patterns (top) and the limited visual detail captured by CLIP features (bottom).
Appendix F Additional Experimental Results
Analysis of captioning operation.
We present additional results from the analysis experiments conducted in Section 3, as shown in Figure 14. These results support our earlier findings that single-pass captioners tend to adopt diffuse attention patterns, and that CLIP’s final-layer features may lack the granularity needed for fine-grained description.
Comparison between our specialist and LLaVA-1.5-7B.
Additional results comparing our lightweight specialist with the large multimodal generalist (LLaVA-1.5-7B), as discussed in Section 3, are shown in Figure 15. Despite its limited parameter size and a frozen vision encoder, our specialist demonstrates unexpectedly strong performance on captioning benchmarks, indicating that small models may still handle detailed captioning tasks reasonably.
Impact of Sharp-Eyed Refinement.
Additional results related to our refinement framework are presented in Figure 16, building on the findings discussed in Section 6. Our refinement method yields caption-quality improvements in both single-sentence and detailed captioning.
Appendix G Extensive Details on The Dataset
G.1 How to generate the pseudo-initial caption $\hat{c}_{k}$?
We describe our strategy to generate the pseudo-initial caption as the following. The instruction prompt used for this strategy is shown in Figure 18.
  • Pseudo-initial captions are generated by modifying 0–3 points from the ground-truth (GT) caption.
  • Modifications are based on four categories: Entity (e.g., chair, cat), Attribute (e.g., color, material like wooden, count like three cups, texture, shape, size, inspired by DSG [1]), Relation (e.g., A in front of B, or B in front of A), Action (e.g., eating, blowing).
  • Except the modifications, the overall sentence structure and style should be preserved.
  • It can occasionally be identical to the GT caption.
  • Few-shot examples are provided like the examples in Table 13.
GT caption pseudo-initial caption
A woman in a room with a cat A woman in a room with two cats
A woman in a room with a dog
A woman in a room with a cat
A multicolored motorcycle rests outside of a sheep farm A multicolored bicycle rests inside a sheep farm
A bright red motorcycle rests outside of a sheep farm
A multicolored motorcycle races around a sheep farm
Table 13: Few-shot examples of GT captions and corresponding pseudo-initial captions.
We consistently generate three pseudo-initial captions per GT caption. Hence, the dataset can be summarized as Table 14, where ‘#pairs’ indicates the number of image-caption pairs.
Dataset #images #GT captions per image #pairs for fine-tuning stage 1 #pseudo-initial captions per GT caption #pairs for fine-tuning stage 2
COCO [14] 113K 5 565K 3 1.6M
ShareGPT4V [11] 100K 5 500K 3 1.5M
DCI [70] 7.4K 10 74K 3 0.2M
GLaMM [61] 550K 1 550K 3 1.6M
Table 14: Statistics of datasets used for fine-tuning. Notably, the number of images in fine-tuning stages 1 and 2 are identical.
Figure 15: Qualitative comparison between our lightweight specialist and the large multimodal generalist LLaVA-1.5-7B using MS-COCO [14] (top) and ShareGPT4V [11] & DCI [70] (bottom). Despite its smaller size and simpler architecture, our model produces competitive descriptions.
Figure 16: Qualitative examples of our sharp-eyed refinement. The results show improved caption quality after refinement, using MS-COCO [14] (top) and ShareGPT4V [11] and DCI [70] (bottom).
G.2 Datasets for Detailed Captioning
While MS COCO [14] has long served as the standard benchmark for image captioning, its annotations are typically limited to single-sentence descriptions that often fail to capture the richness of visual content. Recent works [88, 21, 41, 18] have highlighted this limitation and emphasized the need for more detailed captioning datasets. In this study, we leverage three datasets that offer relatively high-quality and more comprehensive image descriptions: ShareGPT-4V [11], DCI [70], and GLaMM [61]. For ShareGPT-4V, captions are initially generated by GPT-4o and subsequently refined by human annotators. DCI contains human-authored captions. Finally, GLaMM generates detailed captions by aggregating outputs from numerous open-source tools, including object detectors and scene parsers, and composing them using an LLM.
G.3 ShareGPT & DCI
The DCI dataset [70] contains 7.4K training images and 0.4K test images, with each image paired with 10 human-written captions averaging 55 words in length.
The ShareGPT4V dataset [11] includes 100K images, each accompanied by a single long and human-verified caption of approximately 200 words. Using this format directly poses a challenge for fair evaluation with n-gram-based metrics, which benefit from having multiple reference captions per image. So, we summarized the original caption into five captions, each containing roughly 50 words, since generating multiple captions while preserving the original length could introduce hallucinations. The prompt used for this summarization process is presented in the next section, and the resulting dataset will be made publicly available.
Given the relatively small scale of DCI compared to MS COCO (which contains 118K images with five captions per image), we combined DCI with the processed ShareGPT4V dataset to construct a unified benchmark. This yielded 102.4K training images (100K from ShareGPT4V and 7.4K from DCI), each with ten or five detailed captions, and 5K test images—comparable in scale to MS COCO.
G.4 GLaMM
The GLaMM dataset [61] contains automatically generated captions produced using a combination of object detection, scene-graph parsing models, and LLMs. On average, each caption consists of approximately 45 words. In our experience, despite GLaMM’s reliance on extensive visual tools, the quality of the generated captions was often inconsistent. Frequent factual inaccuracies were observed, for example, one to two incorrect words appearing in every two and three captions. Each image in GLaMM is paired with a single caption, which presents challenges for n-gram-based evaluation, as previously discussed. Nevertheless, we include this dataset in our experiments. We randomly sampled 600K image-caption pairs from the full dataset, allocating 30K for testing and using the remaining 570K for training. Possibly due to the quality issues mentioned above, models trained on GLaMM generally underperformed compared to those trained on ShareGPT4V and DCI.
Appendix H Prompt Templates
We share the prompt templates used when interacting with OpenAI’s GPT models throughout our study. First, the prompt used for summarizing long captions into shorter ones, introduced in Section G.3, is shown in Figure 17. Second, the prompt used to generate pseudo initial captions for our refinement framework, as described in Section 5, is provided in Figure 18. The examples generated using it are in Figure 20. Lastly, the prompt used in the MLLM-as-judge evaluation setup is shown in Figure 19. We note that we strictly follow the template proposed in the original MLLM-as-judge papers [8, 9].
Figure 17: Prompt used for summarizing long-form captions into shorter, multi-reference captions.
Figure 18: Prompt used to generate pseudo initial captions for Sharp-Eyed Refinement.
Figure 19: Prompt used for MLLM-as-judge evaluation, following the original template from [8, 9].
Figure 20: Examples of pseudo initial captions generated using the prompt in Figure 18, from MS COCO (top) and ShareGPT4V & DCI (bottom).
Appendix I Experimental Details
I.1 Pretraining and Finetuning
Our implementation is initialized from the LLaVA-1.5 repository [43]. To reduce reliance on LLMs, we replace LLaMA with the OPT series of language models [93]. The training process consists of three stages: pretraining, finetuning for caption generation (in Section 3), and finetuning for sharp-eyed refinement (in Section 5). The hyperparameters used for each stage are summarized in Table 15. We closely follow the original LLaVA training setup. As emphasized in Section E.1, our method differs from previous small-model approaches in that the visual features are directly used as self-attention inputs to the language model.
Pretraining Fine-tuning
Dataset LCS-558K [43] MS COCO [14] or SharedGPT [11] + DCI [70]
Adapter 4-layer MLP with GELU 4-layer MLP with GELU / Deeplens
Trainable Adapter layers only Adapters + Language Model
Training Epochs 1 10 / 2
Learning Rate $1\times 10^{-4}$ $2\times 10^{-5}$
Weight Decay 0 0
Warm-up Ratio 0.03 0.03
Learning Rate Scheduler Cosine decay Cosine decay
Table 15: Hyperparameters used for model training. The settings to the left and right of the / correspond to those used in Section 4 and Section 3, respectively.
I.2 Experiment setting of large multimodal generalist
Table 2 compares specialist models with several generalist MLLMs, including InstructBLIP [15], Unified-IO-XL [47], Shikra [10], Qwen-VL [3], and LLaVA-1.5 [43]. These generalist models were reported to be trained on MS COCO images, but not on datasets such as ShareGPT4V, DCI, or GLaMM. Instead, they were instruction-tuned on large-scale datasets; for example, Qwen-VL and InstructBLIP were trained with approximately 1.5B and 130M instruction samples, respectively. For this reason, we categorize them as generalist models.
The generalist MLLMs were evaluated using publicly available checkpoints from their official GitHub repositories, without any additional fine-tuning. For the single-sentence captioning task, we used the prompt: “Provide a one-sentence caption for the provided image.” For the detailed captioning task, we used the prompt: “Describe the photo within 55 words.”. Additionally, we conduct an experiment on LLaVA-1.5 with the alternative instruction, as shown in Table 16. These results reveal two important points: (1) evaluation metrics such as BERTScore [94] penalize long-form outputs, and (2) longer generations tend to induce more hallucinations, consistent with prior observations [26].
LLaVA-1.5 CIDEr [71] BERTScore [94] CAPT [18]
“Describe the photo within 55 words” 36.1 36.6 48.6
“Describe the photo in detail” 12.8 17.6 40.6
Table 16: Performance of LLaVA-1.5 under different prompt instructions.
I.3 Attention map visualization
As part of our analysis in Section 4, we evaluate whether the model attends to the appropriate image regions when generating specific words in a caption. To this end, we adapt the visualization code provided by API [89], originally designed to point out attention maps between images and questions in VQA tasks. We modify the attention hooking module (https://github.com/yu-rp/apiprompting/blob/master/API/API_LLaVA/hook.py) to visualize attention between image regions and selected words within captions. This allows us to inspect which parts of the image the LM focuses on when producing certain tokens, such as the phrase ‘a white toilet’ in Figure 5.
I.4 Image reconstruction
To investigate whether CLIP’s visual representations are coarse or ambiguous, we conducted an image reconstruction experiment using a Masked AutoEncoder (MAE) framework in Section 4. In this setup, a ViT encoder produces visual features, which the decoder then uses to reconstruct the image. We adopt the same visual encoder as used in LLaVA-1.5, ‘CLIP ViT-L/14-336’, and keep its parameters frozen. The decoder receives the visual embeddings without masking and predicts the corresponding RGB image. We utilize two types of visual inputs: (i) last-layer features (i.e., LF), and (ii) multi-level features (i.e., ML), where ML consists of outputs from layers 13, 18, and 23 of the encoder. The decoder is trained on 100K images from the MS COCO dataset. Further architectural details are available in the MAE repository, https://github.com/facebookresearch/mae/blob/main/models_mae.py. The hyperparameters for optimizing the decoder are summarized in the table below.
Optimizer Learning Rate $\beta$ Weight Decay Batch Size Epoch
AdamW 1e-5 (0.9, 0.95) 0.05 64 10
Table 17: Hyperparameters used for training the MAE decoder in the image reconstruction experiment on MS COCO.
I.5 Long Range Video Question Answering
To assess the feasibility of deploying lightweight captioners in real-world applications, we evaluate our model on the Long-Range Video Question Answering task [91]. This task requires the system to answer multiple-choice questions based on user queries about videos that are 10 minutes or longer in duration. We emphasize that there are currently no video MLLMs capable of handling this task directly. Most existing models [34, 4, 97] impose limits on the number of input video frames they can process, making it difficult to cover the full temporal span of long videos.
To address this, recent approaches [86, 75] propose first extracting per-frame captions and then injecting both the captions and the question into an LLM, which can handle over 100K tokens in a single prompt. For evaluation, we follow the setup provided in the official implementation of LLoVi (https://github.com/CeeZh/LLoVi), replacing only the captioning models. When using the generalist model, we generate a caption for each frame using the prompt: "Describe this frame within 50 words". For our specialist model, we use the same prompt and apply the version fine-tuned on the ShareGPT4V [11] and DCI [70] datasets.